### v1

# An Extensible Software Transport Layer for GPU Networking

Abstract.
Fast-evolving machine learning (ML) workloads have increasing requirements for networking. However, host network transport on RDMA NICs is hard to evolve, causing problems for ML workloads. For example, single-path RDMA traffic is prone to flow collisions that severely degrade collective communication performance. We present UCCL, an extensible software transport layer to evolve GPU networking. UCCL decouples the data path and control path of existing RDMA NICs and efficiently runs the control-path transport on host CPUs. This software extensibility brings in transport innovations that cannot be achieved in hardware for ML workloads, e.g., a multipath transport to resolve flow collisions. ML collectives atop UCCL achieve up to 3.3 higher performance compared to an industry solution.
1. Introduction
Machine learning (ML) workloads and their requirements for networking are evolving rapidly. Less than ten years ago, deep neural networks only had millions of parameters, and were trained atop hundreds of CPUs/GPUs with parameter servers or allreduce collective communication (Li et al., 2014). After five years, large language models (LLMs) began to surge with billions of parameters, and were trained atop thousands of more powerful GPUs with multi-level parallelism and diverse collectives like allreduce, allgather, and reduce-scatter (Grattafiori et al., 2024; Barham et al., 2022). In the recent two years, large-scale LLM serving has become the norm; prefill-decode disaggregation (Zhong et al., 2024), as an efficient serving technique, requires fast peer-to-peer communication. This year, serving Mixture-of-Experts (MoE) models like DeepSeek-V3 (Liu et al., 2024) became very popular, featuring challenging all-to-all communication among hundreds of GPUs.
However, networking techniques especially the host network transport on RDMA NICs are hard to adapt and evolve to better suit the needs of ML workloads. Essentially, hardware changes are time-consuming and take much longer time than software changes. This can lead to a mismatch between the application needs and existing hardware optimizations, which often translates into poor performance. For example, Meta has reported that DCQCN (Zhu et al., 2015)—a popular congestion control (CC) algorithm in datacenters supported by RDMA NICs—does not work well for LLM training workloads with low flow entropy and high traffic burstiness (Gangidi et al., 2024). As a result, Meta decided to disable the CC support in NICs and instead implement traffic scheduling at the application layer. Similarly, DeepSeek disabled the CC when running large-scale all-to-all for serving MoE models (DeepSeek, 2025). However, running a large-scale RDMA network without CC is brittle, as it can lead to deadlocks, head-of-line blocking, and pervasive congestion (Hu et al., 2016; Guo et al., 2016; Mittal et al., 2018; Zhu et al., 2015; Bai et al., 2023; Gao et al., 2021).
In another example, Alibaba has observed severe performance degradation for collective communication during LLM training. This was due to the high level of flow collisions, which in turn was caused by the RDMA NICs supporting only single-flow/path per connection (Qian et al., 2024). To avoid this problem, Alibaba has redesigned the network topology for LLM training using a rail-optimized dual-plane architecture. However, such a redesign is costly to build and maintain. As we will show in §3.2, a software-only solution implementing multipath at the application level would have avoided such topology changes.
In §2.2, we discuss four more examples where ML workloads require adapting RDMA hardware support, including network incast in MoE serving, semi-reliable gradient transmission, efficient loss recovery, and avoiding vendor lock-in. Furthermore, the existing hardware-baked host network transport layer makes it difficult to productize new research proposals (Wang et al., 2024; Le et al., 2024; Bonato et al., 2024; Olteanu et al., 2022) to improve the performance of ML workloads in real-world scenarios.
These examples point to network extensibility as one of the key challenges in datacenter networks. In this paper, we focus on addressing this challenge in the context of RDMA-enabled hosts, by proposing and implementing a software-only extensible transport layer called UCCL (Ultra-CCL). This software-only approach makes it much easier to efficiently support new ML workloads, e.g., CC algorithms that can better support ML traffic than existing DCQCN, and multipathing to mitigate flow collisions. To implement this approach, we need to address two challenges: (1) How to decouple the data and control paths for existing RDMA NICs? The data path handles network data transfers for the GPU, while the control path manages transport control decisions like CC, packet reliability, and multipath load balancing (LB). Only if the control path is decoupled from the data path can we implement it on the CPU. (2) How can we achieve hardware-level performance with the control path running on the CPU? This is challenging given the inter-GPU-server high bandwidth, which can exceed 3.2 Tbps (Amazon Web Services, 2025; Microsoft Azure, 2024).
To address the first challenge, UCCL repurposes the features of existing RDMA NICs to achieve efficient data and control path separation. In particular, UCCL leverages RDMA Unreliable Connection (UC) to bypass hardware-baked CC and packet reliability logic, and uses RDMA immediate data to pass transport control states between sender and receiver CPUs. For some NICs that do not support UC, such as AWS EFA (Services, 2025), UCCL leverages the scatter-gather feature over RDMA Unreliable Datagram (UD) to achieve separation.
To address the second challenge, UCCL leverages an array of techniques tailored to the unique characteristics of ML workloads. On the one hand, UCCL leverages GPUDirect (Corporation, 2025d) to alleviate CPU overhead. On the other hand, UCCL employs control coalescing to make transport control decisions for every 32KB data chunk rather than every packet; this works well as many transport decisions such as CC do not require per-packet reaction, but just per-RTT to avoid overreaction (Li et al., 2019; Google, 2024). As a result, UCCL can handle 400 Gbps unidirectional traffic with a single CPU core, and achieve the same message latency as hardware-based transport. UCCL leverages 256 RDMA QPs (Queue Pairs) per connection to perform multipathing, without worrying about the high QP context swapping overhead highlighted in previous work (Wang et al., 2023); this is because ML workloads feature bulk data transfer with mostly MTU-sized large packets, which effectively amortizes the swapping overhead. UCCL’s software transport is practical and economical in modern GPU servers, which usually have hundreds of powerful CPU cores (in fact, these CPU cores are often heavily underutilized).
UCCL provides an expressive yet easy-to-use interface. To demonstrate the versatility of this interface and the power of UCCL’s extensibility, we use three case studies. First, we implement a multipath transport protocol that mitigates flow collisions by leveraging packet spraying, i.e., randomly sending packets from a single connection across different paths (Dixit et al., 2013). This transport implementation achieves 3.3 higher throughput for collective communication over AWS’ existing solution, SRD (Shalev et al., 2020) on EFA NICs, which translates into up to 1.4 speedup for two ML applications. Second, we implement a receiver-driven protocol EQDS (Olteanu et al., 2022) to handle network incast in MoE-like workloads, achieving 4.9 better message tail latency over InfiniBand built-in transport. Third, we implement selective retransmission (Mathis et al., 1996) for efficient transport loss recovery, and show its superiority over RDMA hardware transport under packet loss. These case studies highlight that UCCL can effectively enable innovations in the transport protocols that are hard to implement in today’s network stack without expensive and time-consuming changes.
2. Background and Motivation
2.1. GPU Networking: from RDMA to Collectives
GPU networking is extremely heterogeneous. At the high level, GPU collective communication libraries like NVIDIA NCCL (NVIDIA Corporation, 2025) and AMD RCCL (AMD, 2025) use RDMA and kernel TCP (non-RDMA) for inter-server networking, with RDMA preferred since it is faster and more efficient. RDMA provides various communication primitives called Queue Pairs (QPs) including Reliable Connection (RC), Unreliable Connection (UC), and Unreliable Datagram (UD):
RC provides one-to-one message semantics (up to 1GB per operation), with NIC hardware handling packet reliability and CC. Some vendors like NVIDIA allow disabling CC.
UC also provides one-to-one message semantics but without NIC hardware logic for packet reliability or CC.
UD provides one-to-many datagram semantics, i.e., under one MTU per operation, without packet reliability or CC.
Some cloud providers build their own RDMA NICs and QPs. For example, AWS EFA NICs replace RC with SRD (Scalable Reliable Datagram) (Shalev et al., 2020) that implements datagram semantics, multipathing, packet reliability, and CC.
To use RDMA NICs, the CPU issues verb operations such as two-sided send/recv and one-sided read/write that transfer data over QPs. Internally, the verb builds a Work Queue Entry (WQE) and performs MMIO-write into the RDMA NIC registers. Upon completion, and depending on whether the verb is two-sided, the RDMA NIC generates a Completion Queue Entry (CQE) for the software to consume. Note that UD supports only two-sided verbs, UC supports all verbs except RDMA read, while RC supports all verbs. RDMA traffic could go through different network fabrics such as RoCE (RDMA over Converged Ethernet) and InfiniBand.
Figure 1 shows an overview of how collective communication uses RDMA. Once the ML application calls a collective like allreduce, the collective library launches a reduction kernel on each participant GPU to handle data reduction/copy. Next, the sender CPU issues an RDMA write over RC QPs to transfer data. The receiver CPU polls a completion flag in its memory, which the sender sets upon write completion. The library manages a set of transport buffers on the GPU memory to buffer the RDMA data, and relies on the GPU kernel to copy data between the transport buffers and application tensor buffers. The GPU kernel also performs reductions (e.g., sum, max) on the transport buffers (from multiple sender GPUs) into the tensor buffers.
2.2. Motivation on Extensibility
Host network transport on RDMA NICs is hard to evolve compared to software applications. This creates problems for fast-evolving ML workloads. We have shown two such examples in §1; below, we give four additional examples to motivate the need for transport layer extensibility.
Receiver-driven CC for incast. Recent MoE serving workloads are prone to network incast problems. In DeepSeek’s online deployment of their V3 model (Liu et al., 2024), each of the 320 GPUs holds a single expert module, where hidden states are exchanged between expert modules across GPUs, i.e., Expert Parallelism (EP). As the request pattern and load change over time, some experts become much hotter than others, receiving more network traffic from other experts leading to network incast issues. DeepSeek reports the hottest expert could receive 10 more load than the average one. This network incast can be better handled by receiver-driven CC (Handley et al., 2017; Olteanu et al., 2022) which controls last-hop congestion—unfortunately, there is no receiver-driven CC on commercial RDMA NICs.
Application-transport codesigns. Codesigning applications and transport behaviors could bring huge performance benefits. For example, recent work MLT (Wang et al., 2024) customizes loss recovery behaviors for ML training to allow semi-reliable transmission based on the gradient importance from applications. Despite achieving great performance improvement, it is not feasible to integrate MLT into existing RDMA NICs even for the latest NVIDIA ConnectX-7 (NVIDIA Corporation, 2024a), due to a lack of enough programmability.
Inefficient loss recovery. RDMA NICs are known to perform poorly under packet loss, especially for old-generation NICs (Skiadopoulos et al., 2024; Mittal et al., 2018; Li et al., 2023; Wang et al., 2023). This is caused by the inefficient go-back-N retransmission logic hardcoded on these NICs due to limited on-chip SRAM constraints. As a result, RDMA deployment normally requires Priority Flow Control (PFC) to achieve a lossless network fabric. However, PFC may lead to deadlocks, head-of-line blocking, and victim flows (Hu et al., 2016; Mittal et al., 2018), and its likelihood is higher as the GPU networking bandwidth keeps increasing (for the reason, we kindly refer to page 4 in (Hu et al., 2016)). If we could extend the transport layer of GPU networking with more efficient selective retransmission, we could better handle packet loss and rely less on PFC (Mittal et al., 2018).
Heterogeneous NICs. Datacenters usually consist of multiple generations and vendors of RDMA NICs due to continuous expansion, cost optimization, and to avoid vendor lock-in. While NVIDIA, Broadcom, and AMD all have 400 Gbps RDMA NICs for ML (NVIDIA Corporation, 2024a; Broadcom, 2024; AMD Pensando, 2024), they come with subtly different control path logic like packet reliability and CC; this heterogeneity reduces achievable bandwidth by 2-33 when communicating between NICs from different generations/vendors, as reported by Alibaba (Li et al., 2023). Instead, if we can extensibly align these NICs’ control path logic in software, we could avoid such a severe performance drop.
2.3. Prior Work on Extensibility
Leveraging SmartNICs. Several recent efforts aim to make RDMA transport programmable by offloading it to the SmartNIC RISC cores, but they have constraints in extensibility and performance. Google Falcon SmartNICs (Google, 2024, 2024) only support programming rate update actions for latency-based Swift CC (Kumar et al., 2020) and path selection decisions with limited paths (Qureshi et al., 2022); similar constraints also apply to firmware updates on hardware RDMA NICs. AMD Pensando SmartNICs (AMD Pensando, 2024) support using P4 language (Bosshart et al., 2014) to program their transport layer, but P4 has limited programmability, e.g., hard to implement efficient loss recovery; it is also unclear if they could support receiver-driven CC. AWS EFA SmartNICs (Services, 2025) implement a proprietary multipath reliable transport SRD (Shalev et al., 2020) using NIC ARM cores. However, they perform poorly on collectives, especially for the connection-intensive all-to-all (see §6.1). We attribute this to the fundamentally limited processing power and cache capacity of these wimpy RISC cores (Phothilimthana et al., 2018; Liu et al., 2019). FPGA-based SmartNICs provide higher performance but have limited extensibility due to hardware resource constraints (Wang et al., 2023).
Leveraging CPUs. A line of work has leveraged host CPUs to make better control decisions in GPU networking. ZeroNIC (Skiadopoulos et al., 2024) modifies the NIC hardware to run the control path of RDMA transport on CPUs, while leaving the data path on the NIC following GPUDirect. In contrast, UCCL aims to be more practical without modifying existing hardware; UCCL by design supports efficient multipathing, compared to ZeroNIC’s single path transport. Flor (Li et al., 2023) leverages RDMA UC to bypass the control path of the RDMA hardware and implements flexible software control on CPUs. Flor targets CPU-based storage applications with 100 Gbps traffic per server, while UCCL targets more network-intensive ML applications that have 3.2+ Tbps traffic per server (Amazon Web Services, 2025; Microsoft Azure, 2024) and develops multipathing to avoid flow collisions. To this end, UCCL adopts several different designs such as multi-QP and connection splitting (see §3.2 and §3.3).
Other efforts design transport protocols to address specific network challenges. Ultra Ethernet Consortium (UEC) (Consortium, 2023) standardizes several multipath transport protocols with packet spraying to address flow collisions for ML workloads: a sender-driven one based on STrack (Le et al., 2024) and SMaRTT-REPS (Bonato et al., 2024), and a receiver-driven one based on EQDS (Olteanu et al., 2022). Prior to UEC, MP-RDMA (Lu et al., 2018) and MPTCP (Multipath TCP community, 2023) designed multipath protocols for CPU workloads to improve robustness under network failures. Overall, these protocols leverage various congestion signals such as Explicit Congestion Notification (ECN) (Alizadeh et al., 2010), RTT (Mittal et al., 2015), and packet trimming status (Cheng et al., 2014; Handley et al., 2017; Olteanu et al., 2022) to make multipath CC and LB decisions.
How to separate control and data (§3.1)? | How to reassemble packets on receivers (§3.1)? | How to enable multipathing (§3.2)? | How to build efficient software transport (§3.3)? | |
RC, UC | Write with immediate | RDMA NIC handles | Different QP connections | Control coalescing |
UD | Send/Recv with scatter-gather | GPU does scattered memcpy | Src-dst QP combinations | Chained posting |
AF_XDP | – | CPU handles | Different UDP ports | User-space packet IO |
3. UCCL Design
Figure 2 shows the high-level architecture of UCCL. UCCL layer sits between the collective library such as NCCL and the low-level communication primitives exposed by NIC hardware, e.g., RC, UC, and UD for RDMA NICs and AF_XDP (The Linux kernel development community, 2025) (a user-space fast packet IO) for non-RDMA NICs. ML applications use collective APIs such as allreduce and point-to-point APIs such as SendRecv exposed by the collective library without directly interacting with the UCCL layer. Both the collective library and the UCCL layer are compiled into individual shared libraries, i.e., libnccl.so and libnccl-net.so for NCCL, providing a drop-in replacement for ML applications without code modification or recompilation. UCCL leverages the network plugin system of existing collective libraries (Corporation, 2025b; AMD, 2025) to avoid changing the library code in most cases, with exceptions for UCCL over UD that requires a slight code modification (detailed in §3.1). For brevity, the remaining paper targets RDMA NICs such as NVIDIA ConnectX NICs and AWS EFA NICs (Services, 2025); we will explicitly mention when targeting non-RDMA NICs.
Figure 2 shows the threading model of the UCCL layer. The UCCL plugin interacts with a group of UCCL engine threads via shared memory to create connections, register/deregister GPU memory regions with the RDMA NICs, and send/recv/flush/poll network messages. Each engine thread runs TX, RX, and pacing functionalities of UCCL multipath reliable transport for multiple UCCL connections. As described in §3.1, UCCL engines instruct the RDMA NIC to receive network data, split control headers and application data payloads, and directly DMA them into the CPU and GPU memory separately. The whole process bypasses the packet reliability and CC logic on RDMA NIC hardware as much as possible by using proper RDMA primitives. After getting control headers in CPU, UCCL engines make transport decisions such as CC, LB, and handling packet loss and reordering. Since these decisions are executed by a normal user-space process on the CPU, instead of on RDMA NIC hardware, they can be easily extended by collective library or ML application developers.
In UCCL, all connections between a specific pair of NICs share the same set of QPs (e.g., 256), including the case of multiple GPUs sharing one NIC. This design fully harnesses the underlying multipath datacenter network while not burning too many QPs (§3.2). UCCL further integrates a bag of techniques to run software transport as efficiently as possible, e.g., control coalescing, connection splitting, chained posting, and more (§3.3). UCCL also supports extending transport for non-RDMA NICs through the AF_XDP user-space packet IO, bypassing conventional kernel TCP stacks. Table 1 summarizes the UCCL techniques to achieve transport extensibility across different communication primitives. We now describe them in the rest of this section.
3.1. Separating Control Path and Data Path
The overall goal of separating the control path and the data path is to enable running extensible transport on the CPU, while efficiently transferring data to/from GPUs in a GPUDirect manner. This goal has three specific aspects: (1) We should involve as little control logic as possible in the data path to let the CPU make more transport decisions, like CC and packet reliability. (2) We must achieve GPUDirect for the data path efficiency (Gershon et al., 2024; Skiadopoulos et al., 2024). (3) We should support heterogeneous RDMA NICs. For example, NVIDIA NICs support UC, while Broadcom and AWS EFA do not.
UC as the preferred QP. UCCL chooses UC as the preferred QP whenever available on the RDMA NIC, similar to Flor (Li et al., 2023), because it supports efficient segmentation and reassembly offloading to the NIC (in contrast to UD), while bypassing hardware-baked CC, loss recovery, and out-of-order packet handling (in contrast to RC). Under UC, UCCL uses the efficient RDMA verb of write with immediate to transfer data chunks; this verb operates in a two-sided mode, so that both sender and receiver CPUs can react to the data transfers. For the data path, the sender CPU specifies the addresses of source and destination data chunks when issuing the verb, and then the sender NIC will automatically segment the source data chunk into MTU-sized packets with packet headers prepended and send them out. When receiving these packets, the receiver NIC will remove the packet headers and reassemble payloads into the contiguous memory region specified by the sender CPU.
For the control path, write with immediate allows carrying a 32-bit imm_data from the sender CPU to the receiver CPU, serving as the control header for UCCL transport. UC guarantees that any successfully arrived chunk will generate a CQE with the imm_data embedded, which is then consumed by the receiver CPU. Within the 32-bit budget, UCCL allocates 8 bits for connection ID and 7 bits for message ID, supporting 256 connections for each pair of NICs on a UCCL engine and 128 in-flight messages for each connection, which is sufficient for collective communication. UCCL then allocates 8 bits to the chunk sequence number (CSN) to identify the position of a chunk in the message being transferred. Another 1 bit is allocated for marking the last chunk of a message. The remaining 8 bits are reserved for more advanced CC such as receiver-driven ones (see §4.2).
RC with CC disabled. In practice, UC is not always supported across different RDMA NIC vendors (Li et al., 2023), e.g., Broadcom (Broadcom, 2024). In these cases, UCCL will opt to RC with CC configured to be disabled, and then leverage RDMA write with immediate in a similar way as UC. On the one hand, RC prevents UCCL from customizing the packet reliability mechanism baked in NIC hardware; on the other hand, it allows faster ACKs and more precise RTT estimation in hardware.
UD as the last resort. Some RDMA NICs do not allow disabling CC for their RC QPs, e.g., AWS EFA NICs (to be precise, EFA NICs do not have RC but only SRD). To support them, UCCL leverages UD at the cost of higher CPU usage compared to UC and RC. On the one hand, UD totally bypasses any hardware control logic on RDMA NICs, well-aligned with our goal. On the other hand, UD only supports sending/receiving MTU-sized data (i.e., no segmentation or reassembly offloading); so UCCL needs to consume more CPU cycles to do segmentation and reassembly.
Challenge of separation. One key challenge for UCCL over UD is that UD does not support RDMA write with immediate (but only the send/recv verbs), so that UCCL over UD cannot designate the immediate data as the transport control header. Then how can UCCL separate control header and data payload using UD (i.e., placing them to CPU and GPU separately)? UCCL must guarantee that the control header and data payload are fate-sharing in terms of loss status and arrival order, so that UCCL can make valid transport decisions based on the control header. One strawman solution would be transferring the control header and data payload together as a single packet into the destination GPU memory, and then the CPU reading the control header from the GPU memory. But this incurs additional performance overhead.
UCCL’s approach is to leverage the scatter-gather feature to let the NIC hardware automatically merge the control header and data payload during RDMA send, and split these two during RDMA recv; Figure 3 shows an example. On the sender side, the CPU issues an RDMA send verb with a two-entry sg_list that specifies the control header address+length on the CPU, and the data payload address+length on the GPU. The RDMA NIC will then read the header and payload from the CPU and GPU respectively, and merge them into a single network packet to send out, as long as the total length does not exceed the MTU size. On the receiver side, the CPU pre-posts a recv verb with a two-entry sg_list that specifies the receiving address+length for the header and payload, respectively. Note that the header length must be a fixed value that the sender and receiver agree on, e.g., 64B in this example; the payload length specified in the recv verb need not exactly match the send verb, but should be no smaller. Later, when the packet arrives at the receiver NIC, the NIC will automatically split the header and payload across CPU and GPU, following the boundary of the fixed header length. UCCL’s approach always keeps the control header fate-sharing with the data payload, and avoids the CPU reading any extra header from the GPU.
Challenge of reassembly. UCCL over UD still faces another challenge of how to correctly and efficiently reassemble packets on the receiver GPU. Recall that UD does not support reassembly offloading to the NIC, and only allows sending/receiving a single packet in one verb (§2.1). We note that sender-side segmentation is relatively easy, as the CPU could partition the transport sending buffer into individual data payloads (based on MTU size), and specify their addresses in send verbs. However, for the receiver-side reassembly, even if the CPU pre-posts recv verbs that specify in-order individual data payload addresses drawn from the transport receiving buffer, the packets will land into the buffer in an out-of-order manner due to packet loss or reordering over the multipath network. This reassembly challenge is unique to UD, as UC/RC allows the sender to directly specify the receiver-side GPU buffer addresses in the write with immediate verb.
Addressing this challenge requires some form of scattered memcpy GPU kernel that copies out-of-order data payloads to the transport receiving buffer (following the right order given by the receiver CPU). But the question is where to launch and run such a kernel. To avoid extra kernel launching overhead, UCCL chooses to fuse such scattered memcpy operations into the existing reduction kernel in collective libraries (§2.1). Our fused kernel will first do scattered memcpy to copy out-of-order data payloads into the transport buffers, then do the original reduction work from the transport buffers to the application tensor buffers. The only overhead of this approach is the extra GPU memory bandwidth consumption, but this is bounded by the network bandwidth. Given the high GPU memory bandwidth (e.g., 1.6-2.0 TB/s in A100), such extra bandwidth consumption is negligible.
For non-RDMA NICs, UCCL builds a reliable transport atop UDP with the AF_XDP technique, an efficient kernel socket that lets the NIC directly DMA network packets to user-space memory regions. We choose AF_XDP as it achieves similar high performance as DPDK (Zhou et al., 2024), but is kernel-native and does not require special NIC drivers, thus easy to deploy (Tu et al., 2021). Similar to how collective libraries such as NCCL use kernel TCP for non-RDMA NICs, UCCL over AF_XDP does packet reassembly on the CPU, followed by a cudaMemcpy() to transfer the received message to the GPU.
3.2. Harnessing Multipath
One of the key motivations for GPU network extensibility is to harness the multipath capacity of modern datacenter networks (§2.2). UCCL achieves this by using multiple UC, RC, or UD QPs. Basically, network traffic from different QPs will likely go through different network paths, as both RoCE and Infiniband usually use ECMP (Equal-Cost Multi-Path) for multipath routing with source and destination QP numbers as the hash inputs (Gangidi et al., 2024; NVIDIA Corporation, 2024b). For UC and RC, UCCL by default uses 256 QPs, which provides maximum 256 different network paths as used by recent transport research (Le et al., 2024; Bonato et al., 2024). For UD, UCCL uses a much smaller number of QPs by combining different source and destination QPs. For example, 16 source UD QPs and 16 destination UD QPs will provide maximum 1616=256 different network paths, because for connection-less UD, each source QP can send packets to any destination QP. UCCL also supports a configurable number of QPs for different collectives, e.g., a smaller number might work well for all-to-all with relatively high entropy (Gangidi et al., 2024). To avoid burning too many QPs, especially when the collective library creates multiple connections between the same pair of NICs (e.g., because of multiple GPUs sharing one NIC), UCCL lets all these connections share the same set of QPs.
We note that making this multi-QP design choice is not trivial, especially as a line of prior work has highlighted the severe QP scalability issues on RDMA NICs (Kalia et al., 2016, 2019; Chen et al., 2019; Monga et al., 2021; Kong et al., 2022; Wang et al., 2023; Lee et al., 2024). For example, SRNIC (Wang et al., 2023) reports 23% bandwidth drop when scaling RC QPs from 256 to 512, and 46% drop when scaling to 16k. This bandwidth drop is caused by the QP swapping overhead: the NIC can only hold/cache limited QP contexts on its SRAM, and must spill/swap the excessive rest to the host DRAM over PCIe, incurring frequent QP swapping. Surprisingly, we do not observe such a severe performance drop for collective communication—we only see 17% drop when scaling RC QPs to 60k (see §6.4.1); for UC with a smaller QP context size, the drop is even negligible.
There are two reasons behind this counter-intuitive phenomenon. First, ML workloads feature large messages, and thus collectives mostly transfer MTU-sized packets; such large transfers effectively amortize the QP swapping overhead. Second, with GPUDirect, GPU data transfers only go through the PCIe switch, but not the PCIe root complex that is connected to the CPU (Kong et al., 2023); therefore, there is no PCIe contention between GPU-NIC traffic and CPU-NIC traffic (that is incurred when NIC swaps QP contexts and CPU posts verbs). In contrast, prior work focuses on small-message CPU workloads such as in-memory Key-Value stores that only transfer dozens of bytes at once for each QP, thus bottlenecked by the QP swapping overhead and PCIe traffic contention. Later in §3.3, we show that UCCL optimizes transport efficiency by transferring data in a larger chunk granularity (e.g., 32KB), further reducing the QP swapping overhead.
For non-RDMA NIC multipathing, UCCL specifies different UDP ports in packets before sending them out in AF_XDP. This adds no overhead compared to single-path transport.
Handling out-of-order packets. Many factors could cause out-of-order packet delivery, including multipathing, packet loss, and the unpredictable multi-QP scheduler in RDMA hardware (Wang et al., 2023). Existing RDMA NICs perform poorly when handling out-of-order packets, as they cannot maintain large reordering buffers and states due to limited on-chip SRAM constraints (Mittal et al., 2018; Wang et al., 2023). In contrast, UCCL is able to handle out-of-order packets efficiently thanks to its software flexibility and separation of data and control paths. Basically, UCCL follows typical TCP designs with seq and ack numbers to guide packet reordering, fast retransmission (upon duplicate ACKs), and timeout retransmission. UCCL sets a larger duplicate ACK threshold for fast retransmission, instead of the default three in TCP, to accommodate more frequent packet reordering caused by multipathing. Different from TCP, UCCL maintains its packet reordering buffers in the GPU memory and lets the NIC directly DMA network data there. Figure 4 depicts this process with examples. For UC/RC, the reordering buffers are individual data chunks, and the sender CPU specifies in-order chunk addresses when posting verbs. For UD, the reordering buffers are individual packet payloads, and the GPU reduction kernel reorders packets when copying them into the transport buffers (§3.1).
3.3. Towards Efficient Software Transport
So far, we have discussed how UCCL decouples the control path and data path to make flexible transport decisions on the CPU, and how UCCL achieves multipathing. The next question is how to efficiently implement a software multipath transport to support the high bandwidth in GPU networking. This is challenging as a single GPU server could have 8400 Gbps RDMA NICs, totaling 3.2 Tbps bandwidth bidirectionally (Huang et al., 2012); the next generation RDMA NIC will achieve 800 Gbps (Corporation, 2025c), rendering 6.4 Tbps bandwidth. As a reference, Google’s software transport Snap (Marty et al., 2019) can handle 80 Gbps traffic on a CPU core (though they do not use RDMA NICs). Our goal is to use 1 CPU core to handle 400G unidirectional traffic. To this end, we leverage the following techniques:
Run-to-completion execution. Each UCCL engine thread runs RX, TX, pacing, timeout detection, and retransmission functionalities for a set of connections in an efficient run-to-completion manner (Belay et al., 2014; Jamshed et al., 2017). UCCL employs Deficit Round Robin (DRR) (Shreedhar and Varghese, 1995) scheduling to fairly multiplex one engine thread among multiple functionalities and connections.
Connection splitting. To handle 400+ Gbps traffic per NIC more efficiently, UCCL pivots away from the Flor (Li et al., 2023) design of a single CPU core for one connection, but leverages multiple cores for one connection with connection splitting. Basically, UCCL equally partitions the 256 QPs among all engine threads responsible for a specific NIC; each engine thread gets its own connection states for CC and LB, forming a sub-connection. Within each sub-connection, UCCL uses RDMA SRQ and SCQ (Shared Recv/Completion Queues) to reduce the overhead when polling multiple recv and completion queues. The application threads atop the UCCL plugin are responsible for choosing the least-loaded engine (e.g., the engine with the least unconsumed messages) when dispatching messages via SHM. In this way, UCCL could scale transport processing of a single connection to multiple cores, and handle transient load imbalance among CPUs at runtime. It also reduces TX packet bursts by avoiding sending all messages at once from a single core.
Control coalescing. There is an inherent tradeoff between the control decision granularity and software transport efficiency. One could run CC, LB, and reliability logic for each packet to achieve precise control of the transport behaviors, at the cost of consuming more CPU cores. Alternatively, one could relax the control granularity by coalescing several same-path packets and making control decisions together, thus with lower CPU consumption. For UC/RC, this also means an RDMA write could directly transmit several packets as a single data chunk, leveraging NIC-offloaded segmentation and reassembly. UCCL employs this control coalescing design with 32KB chunk size as default, striking a balanced tradeoff. Under this chunk size, UCCL can saturate 400 Gbps unidirectional bandwidth with 1 CPU core (§6.4.3), while not severely disrupting transport behaviors/performance (see our packet-level simulation in §C.4.1). Nevertheless, UCCL could also adaptively adjust chunk size based on the congestion level, e.g., switching to a small chunk size to make more precise control when congestion window (cwnd) drops below a threshold or severe packet loss happens.
Chained posting. UD does not support NIC offloading for segmentation and reassembly, thus it incurs more MMIO writes than UC/RC when issuing send/recv verbs (e.g., for individual packets). To reduce such overhead, UCCL leverages the chained posting feature of RDMA NICs to issue one MMIO write for posting up to 32 send/recv verbs. Concretely, the WQEs of these 32 verbs are chained together through the next pointer in previous WQEs, and get posted to the RDMA NIC in one MMIO write.
Small message bypassing multipath. Using 256 UC/RC QPs for small messages still incurs a high performance penalty in terms of latency; this is because these small message transfers cannot fully amortize the QP swapping overhead. To optimize this, UCCL empirically bypasses multipathing for messages under 8KB and transmits them directly via a dedicated UC/RC QP. We argue this bypassing design for small messages does not impact the overall benefits of multipathing, which usually begins to matter at high traffic load with large message transfers (Le et al., 2024).
3.4. Congestion Signals
Relatively restricted congestion signal is a common limitation for software-based reliable transport atop RDMA NICs, including UCCL and Flor (Li et al., 2023). This is because existing RDMA NICs consume packet headers that contain congestion signals like ECN marks (Alizadeh et al., 2010) and packet trimming status (Cheng et al., 2014; Handley et al., 2017; Olteanu et al., 2022), and deliver only the packet payload to the software. Fortunately, the software can still use the RTT congestion signal by leveraging hardware TX/RX timestamping supported by many RDMA NICs, and rely on packet loss as the last-resort congestion signal. Therefore, our current UCCL implementation uses per-path RTT and packet loss to detect congestion and choose paths. In fact, latency-based CC and LB are being widely used in Google’s datacenters (Kumar et al., 2020; Qureshi et al., 2022).
Signal fidelity. Running CC and LB in software using RTT also raises signal fidelity concerns. Overall, there are three factors affecting the fidelity: accurate RTT estimation at the sender, CC decision delay at either the sender or receiver, i.e., the software delay from receiving the congestion signal (e.g., ACK-derived timestamps) to updating the congestion window/rate, and ACK turnaround delay at the receiver, i.e., delay between receiving data chunks and sending back ACK. For , UCCL leverages the NIC hardware timestamps and excludes ACK turnaround delay from the RTT (similar to Swift (Kumar et al., 2020)). For and , theoretically, these delays impact how fast the sender can react to network condition changes, thus impacting the decision precision; however, in practice, even hardware-based transport handles CC events in a per-RTT granularity (e.g., tens of microseconds) rather than per-ACK to avoid overreaction (Li et al., 2019). For example, Google Falcon hardware transport runs CC to update rate once per RTT (Google, 2024). Thus, the few microseconds of decision delay introduced by the software is negligible. Nevertheless, UCCL still employs several techniques to reduce the two delays: similar to Flor (Li et al., 2023), UCCL uses a dedicated high-priority QP for ACK (using in-network priority such as DSCP) and always first polls its completion queue; UCCL further allocates the ACK polling a higher processing budget during DRR scheduling. We quantify these delays in §6.5.2.
4. Extensibility Case Studies
UCCL provides expressive interfaces to implement and extend multipath transport. Due to space limitations, we elaborate on them in Appendix A. Collective library and application developers could also directly extend UCCL transport code, e.g., with the new loss recovery scheme in MLT (Wang et al., 2024), and deploy it quickly in a normal user-space process. We now demonstrate how UCCL extensibility enables new transport designs that work best for different ML workloads.
4.1. Multipath Transport with Packet Spraying
Recent transport research (Le et al., 2024; Bonato et al., 2024) and UEC advocate packet spraying with hundreds of paths as an effective way to address flow collisions in ML workloads. It is naturally challenging for hardware NICs to implement packet spraying because of excessive per-path states, such as path RTTs. Instead, UCCL can easily support packet spraying by maintaining per-path RTT in software. UCCL’s software transport uses Power-of-Two sampling (Mitzenmacher, 2001) to select a path with the lowest RTT, then runs CC to decide how many packets and what rate to transmit. UCCL implements two CC algorithms: one is CUBIC (Ha et al., 2008) used in Linux kernel TCP as the default CC, and another is Swift (Kumar et al., 2020), an RTT-based CC used by Google. UCCL supports both per-path CC states (e.g., per-path cwnd) and global CC states controlling all paths; both achieve similar collective performance in our testbeds. UCCL by default uses global CC during evaluation.
4.2. Receiver-Driven CC
Receiver-driven transports, such as EQDS (Olteanu et al., 2022), NDP (Handley et al., 2017), and Homa (Montazeri et al., 2018), proactively control packet sending rate at the receiver by allocating credits to senders. These transports have proven to resolve the last-hop congestion for network incast effectively, which could happen in MoE serving (§2.2). However, to the best of our knowledge, there are no off-the-shelf NICs that support receiver-driven transports. One reason is that they are vastly different from popular sender-driven ones, and implementing them would require NIC hardware modification. Instead, UCCL’s extensibility enables developers to quickly implement and tune receiver-driven transports in software. We choose to implement EQDS (Olteanu et al., 2022) in UCCL as an example, the state-of-the-art receiver-driven transport adopted by UEC (Consortium, 2023). Our EQDS implementation closely follows the EQDS paper (Olteanu et al., 2022) with a dedicated pacer thread per NIC on the receiver to issue credit packets for senders. For more details, we refer to Appendix B.
4.3. Efficient Loss Recovery
UCCL allows customizing the transport loss recovery logic to support more advanced mechanisms other than the go-back-N retransmission baked into many RDMA NICs. Go-back-N directly drops out-of-order packets to avoid buffering them on expensive on-chip SRAM, but this gives poor performance when packet loss happens (Kalia et al., 2019; Wang et al., 2023; Zhu et al., 2015). Instead, we implement a more efficient selective retransmission (Mathis et al., 1996) in UCCL, by maintaining reordering buffers in the GPU memory (§3.2). Our implementation follows the standard selective retransmission mechanism in TCP, and uses a std::map to track an arbitrary number of out-of-order packets. With more efficient loss recovery in UCCL, ML workloads could possibly run in lossy datacenter networks without PFC (§2.2).
We note that the latest generation of NVIDIA RDMA NICs (NVIDIA Corporation, 2024a) has implemented a limited form of selective retransmission with a small fixed tracking window (to keep low SRAM usage). It will fall back to go-back-N when the number of unacked inflight packets goes beyond that window; in ML clusters with increasing networking bandwidth and thus more inflight packets, such hardware-based loss recovery can hardly work as effectively as the flexible software-based one in UCCL.
5. Implementation
We implement UCCL in 27.4K lines of C++ as a network plugin to NCCL, the standard collective library for NVIDIA GPUs. Our current implementation supports NVIDIA RDMA NICs via RC and UC, AWS EFA NICs via UD, and AWS ENA non-RDMA NICs via AF_XDP. UCCL leverages the popular libibverbs for RDMA NICs and kernel built-in AF_XDP for non-RDMA NICs; therefore, it should naturally support other vendors’ NICs. To enable scattered memcpy when using UD, we add or modify 170 LoC to the NCCL codebase, including a scattered memcpy GPU kernel and C++ code that passes packet pointers to the kernel via CPU-GPU shared memory (created by cudaHostAlloc()). We added two new interfaces to the NCCL network plugin system: irecv_scattered that receives data in a list of scattered packets, and irecv_free_ptrs that frees the scattered packet buffers after scattered memcpy. UCCL also supports running NCCL with multiple processes.
Name | # of Servers | Network | Topology | NIC | GPU | CPU |
CX | 2 | InfiniBand | Same rack | NVIDIA ConnectX-7 400G 8 | H100-80G 8 | Intel Xeon Platinum 8462Y+ (128 cores) |
EFA | 4 | Ethernet | Fat-tree cross racks | AWS EFA 100G 4 | A100-40G 8 | Intel Xeon Platinum 8575CL (96 cores) |
There are some implementation details worthy of note. For NVIDIA NICs, to calculate network RTT based on NIC hardware TX/RX timestamps and CPU ACK timestamps, we implement a NIC-CPU clock synchronization scheme following Swift (Kumar et al., 2020). We use a synchronization time interval of 100. For AWS EFA NICs that do not support hardware timestamps yet, we calculate RTT based on software timestamps minus estimated packet queueing delay at the host. We estimate the queueing delay by dividing the size of queued packets by the NIC bandwidth. For AWS EFA NICs, we find that allocating UD QPs with consecutive QP numbers for a connection yields much better performance than not. We suspect this is because the EFA NICs do round-robin assignments when mapping a QP (and its associated packets) to a ARM core; consecutive allocations help map the QPs of a connection to different ARM cores, avoiding load imbalance.
6. Evaluation
In this section, we aim to answer the following questions:
What is the collective performance of UCCL software multipath transport compared to hardware-based ones (§6.1)?
Can UCCL improve ML application performance (§6.2)?
Can UCCL extensibility benefit certain workloads (§6.3)?
What is the scalability of UCCL (§6.4)?
How do different designs impact UCCL (§6.5)?
Testbed. Table 2 describes our evaluation testbeds. The CX testbed under the same rack mainly serves as a performance stress-testing for UCCL software implementation, as it has no network congestion or flow collisions. We failed to rent a larger testbed from public cloud providers because they have strict GPU quota limits for individuals. Instead, we try to simulate a larger testbed by disabling NCCL NVLink and SHM communication for some tests. This forces GPUs within a server to communicate through the network, and each GPU essentially behaves like a virtual server, e.g., CX becomes 16 virtual GPU servers each with 400G bandwidth.
Experiment setup. For collective performance evaluation, we use NCCL v2.23.4 (Corporation, 2024b) and NCCL-tests 9d26b84 (Corporation, 2024a), the latest versions when we start integrating UCCL into NCCL, and focus on the representative allreduce and all-to-all collectives. NCCL-tests vary the collective message size and measure the achieved latency and bus bandwidth; the latency is inversely proportional to the bandwidth.
Comparison baselines. On the CX testbed, we compare UCCL to the NCCL built-in RDMA support that uses CX-7 RC. UCCL uses 256 UC/RC QPs for each NIC pair, while the NCCL built-in uses QP scaling with 4 RC QPs per connection (Gangidi et al., 2024; Le et al., 2024). UCCL uses CUBIC CC on CX unless specified, which performs better than Swift in the lossless InfiniBand. Since InfiniBand has no packet loss for CUBIC to reduce cwnd, to avoid severe network congestion, we enforce a maximum cwnd value in UCCL CUBIC CC to limit the maximum inflight bytes, similar to TCP flow control.
On the EFA testbed, we compare UCCL to the official AWS NCCL-EFA plugin v1.13.2 (Amazon Web Services, 2024) that uses the multipath SRD (Shalev et al., 2020). UCCL uses 1026=260 UD QPs to serve all connections, which yields the best performance empirically, while the NCCL-EFA plugin uses the best parameters recommended by AWS. UCCL uses CUBIC CC on EFA by default, as EFA NICs do not support hardware timestamping that is critical to Swift CC. In both testbeds, UCCL uses per-path RTT for multipath LB (§4.1).
As the cost of extensible software transport, NCCL atop UCCL uses more CPU cores than vanilla NCCL. By default, the vanilla NCCL uses 2 cores per GPU, while UCCL uses 2 more cores per NIC to run the engine threads (3 more cores per NIC for receiver-driven CC due to the pacer). For example, on CX, they use 16 and 32 cores respectively; on EFA, 16 and 24 respectively.
6.1. Collective Performance
On CX. Figure 5 compares the allreduce and all-to-all performance of UCCL vs. ConnextX-7 on the CX testbed. UCCL UC/RC performs almost the same as ConnectX-7 with various data sizes, except that UCCL UC performs slightly worse (i.e., less than 4%) than ConnextX-7 and UCCL RC for allreduce after 128MB. For allreduce, this exception is because of the extra overhead of handling packet reliability in UC; for all-to-all, ConnextX-7 and UCCL RC do not perform better than UCCL UC, because all-to-all is connection-intensive, causing more QP swapping for RC. UCCL exhibits lower latency for small messages than ConnectX-7 since UCCL adaptively bypasses multipath (§3.3), while ConnectX-7 always uses 4 RC QPs per connection. These experiments confirm that UCCL can make highly efficient control decisions in software, reaching the performance of ASIC-based RDMA NICs (i.e., NVIDIA ConnectX-7) for ML collectives.
On EFA. Figure 6 compares the collective performance of UCCL and SRD on the EFA testbed. UCCL CUBIC and EQDS achieve similar performance with significant improvement over SRD (except for 256MB allreduce): for allreduce, UCCL outperforms SRD by up to 1.37 and 1.27 for small and large message ranges; for all-to-all, the speedup is up to 2.93 and 3.27 respectively. UCCL outperforms SRD because beefy CPU cores are faster in making transport decisions than wimpy ARM cores on EFA NICs (that run SRD), especially when handling connection-intensive all-to-all. Surprisingly, simple CUBIC CC performs very well. This is mainly because UCCL leverages hundreds of network paths and avoids congested ones before sending packets; so most of the time, CC does not get involved. The small performance gap after 256MB is caused by the extra control header added by UCCL, while SRD directly reuses the UDP header as its control header (UCCL cannot reuse as it is not exposed to CPU software).
Figure 7 shows the collective performance comparison on the EFA testbed when NVLink and SHM are enabled (meaning a smaller testbed). Even though a major amount of data traffic goes through the high-bandwidth NVLink, UCCL still achieves much higher or comparable collective performance than SRD, i.e., up to 1.57 for allreduce and 2.14 for all-to-all. We note that all-to-all with NVLink on performs worse than NVLink off (Figure 6). This is because our current UCCL implementation does not support the NCCL PXN (NVIDIA Corporation, 2022) optimization when NVLink is on. Basically, PXN aggregates network messages destined to the same server, and uses the destination NVLink to segment them into different GPUs, which vastly improves small message latency. However, it involves a complicated multireceive mechanism that causes engineering difficulties for UCCL UD integration (that requires changing the NCCL codebase), which we leave as future development. Nevertheless, UCCL still performs better than SRD under small messages in this experiment.
6.2. Application Performance
We run two applications to evaluate how UCCL improves ML workload performance on the EFA testbed. One is a classic ResNet (He et al., 2016) distributed training in PyTorch with the widely-used 1F1B mechanism (Harlap et al., 2018; Zhang et al., 2017) that tries to overlap communication with computation. However, as the GPU compute capability grows much faster than the network bandwidth, the communication time cannot be entirely hidden by the computation time, especially on EFA. Another is a DeepSeek-V3-like MoE serving application. However, we fail to directly serve DeepSeek-V3 on our testbed, because (1) it uses a customized communication library rather than the standard NCCL, so it cannot directly use UCCL, and (2) our testbed does not have enough GPU memory to fit all model weights. Instead, we get a realistic DeepSeek-V3 trace including GPU compute time and network message sizes (i.e., hidden state sizes) from (zartbot/shallowsim, 2025), and emulate the computation and communication behaviors with PyTorch and NCCL.
Figure 8 shows the results for ResNet distributed training. UCCL reduces the training epoch time by 1.07-1.11 compared to SRD. Figure 8 shows the results for DeepSeek-V3 serving. UCCL reduces the per-request prefilling and decoding latency by 1.13 and 1.42 respectively. These experiments demonstrate that UCCL is able to bring significant benefits to real-world ML applications.
6.3. UCCL Extensibility
6.3.1. Handling network incast
In a shared RDMA network, network incast can lead to victim flows due to PFC propagation (Wang et al., 2023; Hu et al., 2016; Mittal et al., 2018). Such network incast can be caused by the Expert Parallelism in MoE serving (§2.2) and gather/reduce collectives in multi-level parallelism of LLM training (Barham et al., 2022; Grattafiori et al., 2024). This experiment tries to create a scenario where network incast co-exists with other collective traffic, and evaluates how UCCL performs compared to the RDMA hardware transport on CX. We co-locate 15-to-1 incast traffic and 16-NIC permutation traffic (Raiciu et al., 2011; Cao et al., 2013; Le et al., 2024; Bonato et al., 2024), where each NIC streams data to another and no NIC receives more than one stream, representing typical collectives. For both types of traffic, each NIC sends 1MB messages with at most four in flight. We compare the receiver-driven EQDS in UCCL vs. the sender-driven InfiniBand CC in ConnectX-7.
Figure 9 and 9 show the FCT distributions for incast traffic and permutation traffic, respectively. Compared to InfiniBand, UCCL EQDS reduces P99/P99.9 latency by 1.73/1.72 for incast traffic and 4.50/4.88 for permutation traffic. This is because the InfiniBand CC only reduces the sending rate when severe congestion and queue build-up have occurred on the incast switch port; but this is too late, as the Credit-Based Flow Control (InfiniBand Trade Association, 2020) (PFC equivalent in InfiniBand) has already paused all upstream ports and heavily disturbed victim flows (i.e., permutation traffic). Conversely, UCCL EQDS proactively controls the rate of all senders on the receiver side, which reduces queue build-up and avoids upstream ports entering the pausing state.
6.3.2. Handling packet loss
We then evaluate how UCCL selective retransmission handles packet loss, compared to hardware-based loss recovery. Furthermore, UCCL coalesces transport decisions like loss recovery at chunk granularity for high software efficiency; therefore, we also test whether this design choice would cause poor loss recovery performance. To this end, we instrument different packet loss rates in software when running NCCL collectives over UCCL UC. Unfortunately, we cannot instrument packet loss for our CX NICs as that would require reconfiguring the NIC or switch behaviors, both of which we do not have access to. Instead, we cite comparable numbers from prior literature.
Figure 10 shows the results. UCCL has only 3-23%/22-41% performance drop for allreduce and all-to-all under 0.01% packet loss rate, compared to the reported 77% drop of RDMA hardware transport in Flor (Li et al., 2023) (combining Flor Figure 6(a) and 7(a) under 1/16384 loss rate). With 1% high packet loss rate, UCCL still retains 58%/72% of the performance for 1GB messages, compared to old generation RDMA NICs that achieve near-zero throughput as reported in DCQCN (Zhu et al., 2015) (Figure 21). This confirms the effectiveness and advantage of UCCL loss recovery despite control coalescing.
6.4. UCCL Scalabality
6.4.1. Varying number of QPs
We vary the number of QPs per NIC pair from 256 to 4096; given that 1 NIC connects to 15 other NICs in all-to-all, this results in 3840 (3.75K) to 61440 (60K) QPs per NIC. With 60K QPs per NIC, we essentially model the QP swapping overhead for UCCL all-to-all across 241 GPUs using 256 paths per connection (i.e., 61440/256=240, plus 1 to include itself) or 961 GPUs using 64 paths. This has covered the largest-scale collective as far as we know, i.e., 256 GPUs reported by Meta (Gangidi et al., 2024) (DeepSeek-V3 limits each token to be dispatched to 4 servers, though with 320 GPUs in total (Liu et al., 2024)).
Figure 11 shows the results. For RC, using 60K QPs incurs a maximum 17% performance drop compared to 3.75K QPs (that have similar performance as the vanilla NCCL). For UC, the performance remains nearly unchanged across different numbers of QPs. As mentioned in §3.2, this is because UCCL’s chunked data transfers (i.e., control coalescing) for large ML messages effectively amortize the QP swapping overhead while GPUDirect avoids PCIe traffic contention. Note that UCCL on EFA does not have QP scalability issues, as UCCL uses connection-less UD there.
6.4.2. Varying chunk size
Figure 12 shows the all-to-all performance on the CX testbed when varying chunk sizes. Small message performance mostly remains unaffected by different chunk sizes, thus omitted here. Our results reveal that large messages do require a larger chunk size to achieve line rate, i.e., 16KB or 4 MTU-sized chunks. This experiment demonstrates the performance benefits of control coalescing for UCCL software transport.
6.4.3. Varying number of CPU cores
Figure 12 shows the all-to-all performance when varying the number of UCCL engines/CPU cores per NIC. Small message latency exhibits negligible difference across different CPU cores, thus omitted here. For ASIC-based NICs with segmentation and reassembly offloading, even though one CPU core can only saturate bidirectional 29GB/s, two CPU cores are enough to saturate the full 50GB/s line rate. This experiment confirms the high efficiency of UCCL software transport, i.e., 1 CPU core saturating 400G unidirectional traffic.
We show more scalability studies in Appendix C.3 when varying numbers of CPU cores, GPUs, and paths on EFA.
6.5. Design Drill-Down
6.5.1. Connection splitting
We now evaluate how connection splitting impacts collective performance in UCCL. Without connection splitting, each UCCL connection is mapped to a single engine out of the two UCCL engines per NIC; with splitting, each UCCL connection dispatches and load balances its messages among the two UCCL engines. Our experiment results show that UCCL without connection splitting only achieves 45.7 and 39.9 GB/s max busbw for allreduce and all-to-all on CX respectively, while UCCL with splitting achieves 48.9 and 48.5 GB/s respectively, saturating the line rate. This is because connection splitting enables multiple CPU cores to process inflight messages in a load-balanced manner. We anticipate connecting splitting will become more critical as the NIC scales to 800 Gbps (Corporation, 2025c).
6.5.2. Congestion signal fidelity in software
UCCL makes a fundamental shift of moving transport from hardware to software. This experiment evaluates how much software transport sacrifices in terms of congestion signal fidelity. We look at the two important metrics as mentioned in §3.4: CC decision delay and ACK turnaround delay. Table 3 shows the results. At light loads, the two delays are limited to at most 10 in both P50 and P99; at high loads, they grow up to 36 for P99. Both are on par with typical datacenter RTTs of 10-40 (Gibson et al., 2022). Therefore, this experiment confirms that UCCL software transport has high enough CC fidelity to make precise per-RTT transport decisions (Li et al., 2019; Google, 2024).
Delay () | Allreduce Light | All-to-All Light | Allreduce Heavy | All-to-All Heavy |
CC decision P50 | 1.7 | 1.7 | 2.1 | 2.9 |
CC decision P99 | 3.6 | 3.4 | 7.0 | 10.8 |
ACK turnaround P50 | 2.0 | 3.0 | 4.0 | 7.0 |
ACK turnaround P99 | 3.0 | 10.0 | 10.0 | 36.0 |
We show more design drill-downs in Appendix C.4, including the impact of chunk size, LB policy, kernel fusion (for scattered memcpy), and PCIe overhead analyses.
7. Discussion
UCCL works around many challenges to support extensible transport with efficient multipathing on existing RDMA NICs, at the cost of QP swapping overhead (though not high for ML collectives), control coalescing, and more. UCCL performance and control granularity would get further boosted if there were a better HW-SW interface in RDMA NICs. We would like to highlight a few points based on our experience of developing UCCL: (1) UC abstraction is powerful with segmentation and reassembly offloading, but using many UC QPs for multipathing incurs some degree of QP swapping overhead; instead, it should evolve into a multipath UC abstraction that allows the software to specify different flow entropies for different verbs over a single QP, just like how UCCL specifies different UDP ports for AF_XDP. (2) NIC HW should expose more congestion signals to SW, such as ECN marks and packet trimming information; these signals can be embedded into CQEs just like hardware timestamps.
8. Other Related Work
Recent work SCR (Zhao et al., 2025) implements receiver-driven CC and multipathing on the DPA (DataPath Accelerator, essentially 16 RISC-V cores) of NVIDIA BlueField-3 SmartNICs. It would have similar performance issues as AWS EFA NICs. Moreover, DPA programmability is limited to what the NIC hardware supports, e.g., only rate-based control, while packet reliability and retransmission are still baked into the hardware. Because of this, SCR needs to alter the original receiver-driven CC to let credits represent available bandwidth rather than bytes. For multipathing, SCR only demonstrates two paths; given the limited L1/L2 cache in DPA (Zhao et al., 2025), it is not clear if SCR could scale to hundreds of paths (Le et al., 2024; Bonato et al., 2024; Olteanu et al., 2022).
Google GPUDirect-TCPX (Almasry and Willem de Bruijn, Eric Dumazet, Kaiyuan Zhang, 2023) integrates GPUDirect into the kernel TCP stack by leveraging the Header-Data Split feature in certain non-RDMA NICs. Instead, UCCL targets both RDMA and non-RDMA NICs, and further supports efficient multipathing. C4 (Dong et al., 2024) does coarse-grained flow-level traffic planning and path selection in LLM training, without programmability for low-level RDMA transport components such as CC and loss recovery. MSCCL (Azure/msccl, 2024) supports customizing collective communication algorithms and could work together with UCCL. Lastly, UCCL is inspired by a line of work that targets extensibility for CPU applications, e.g., Google 1RMA (Singhvi et al., 2020), eRPC (Kalia et al., 2019), and RoGUE (Le et al., 2018) for extensible RDMA transport, and SPIN (Bershad et al., 1995), Exokernel (Engler et al., 1995), and VINO (Seltzer et al., 1996) for extensible OS.
9. Conclusion
UCCL is an extensible and efficient software transport layer for GPU networking. It achieves network extensibility by separating the control path and data path for existing RDMA NICs and running the transport control path in software. Meanwhile, it achieves hardware-level performance by leveraging techniques like control coalescing and connection splitting. We hope UCCL opens the door to the productization of new research proposals on network transports for ML workloads. We will open source UCCL.
Appendix A Interface to Extensibility
By executing control decisions in software, UCCL allows flexibly extending its transport implementation for different scenarios. To ease the development of new multipath transport such as new CC or LB policies among different paths, UCCL exposes a set of expressive interfaces to collective library or ML application developers, as shown in Listing 1.
onChunkSize is called when UCCL chunks a message for transmission, and it returns the permitted chunk size for now. CC could enforce window control here. After it returns, UCCL will build a chunk_desc for the chunk.
onPacingChunk determines if a chunk needs to queue in the timing wheel for rate pacing, and returns true if so.
onSelectPath is called when a chunk is ready for transmission. conn_state contains rich information for selecting path, e.g., RTT scoreboard for each path. It returns the selected path_id (i.e., QP ID) for transmission.
onTxRtxChunk is called when the reliable transport wants to retransmit a chunk, and it returns true if this chunk is permitted to retransmit. CC could enforce window control for the retransmitted chunk here.
onRxChunk is called when receiving a data chunk.
onRxRtxChunk is called when receiving a retransmitted chunk. CC could react to the retransmission chunk here.
onRxACK is called when receiving an ACK. CC could react to the ACK here.
onRxCredit is called when receiving a credit. Receiver-driven CC could react to the credit here.
Appendix B EQDS Implementation Under UCCL
Figure 13 shows the overall implementation. For each NIC, UCCL creates a dedicated pacer thread that operates at a constant rate (derived from the NIC bandwidth) to select candidate senders, allocate credits, and send credits following the EQDS algorithm. Each pacer thread uses a credit UD QP for sending credit packets, and each TX&RX thread also has its own credit QP for receiving credit packets from remote pacers. The pacer thread maintains three lists, i.e., rtx (retransmission), active, and idle sender list, with priorities from high to low. After the TX&RX threads receive data chunks, they will notify the pacer thread via an efficient atomic write in the SHM. Then the pacer will update the sender list: senders who encounter packet loss will be put into the rtx list; senders who have satisfied their requirements will be put into the idle list; otherwise, they will be put into the active list. One note is that as packet tramming is not available in our RDMA NICs and switches, UCCL uses timeout + RTS (Request-To-Send) as a replacement, as suggested by the EQDS paper.
Appendix C More Evaluation Results
C.1. More collectives in ML workloads
C.1.1. Allgather and reduce-scatter
These two are representative collectives used in PyTorch FSDP (Fully Sharded Data Parallelism) (Zhao et al., 2023), exhibiting low network congestion (Gangidi et al., 2024). Figure 14 and Figure 14 compare the allgather and reduce-scatter performance of UCCL and SRD on the EFA testbed. UCCL outperforms SRD by up to 1.68 for allgather and up to 2.18 for reduce-scatter.
C.1.2. Multi-collectives
In multi-collectives, GPUs with the same local rank in each server form a collective group, and multiple groups conduct collectives in parallel. For example, multi-allreduce is used in ML workloads with intra-server Tensor Parallelism (TP) + inter-server Data Parallelism (DP). We evaluate multi-collective performance by setting the environment variable NCCL_TESTS_SPLIT_MASK=0x7 in NCCL-tests. Figure 15, 15, 15 and 15 compare the multi-allreduce, multi-all-to-all, multi-allgather and multi-reduce-scatter performance of UCCL and SRD on the EFA testbed. UCCL outperforms SRD by up to 1.54, 1.22, 1.46, and 1.44 for multi-allreduce, multi-all-to-all, multi-allgather and multi-reduce-scatter, respectively.
C.2. UCCL AF_XDP performance
Figure 16 compares the collective performance of UCCL AF_XDP vs. NCCL kernel TCP with and without proximity placement group (PPG). This experiment is done on AWS with two 50 Gbps VMs connected through AWS ENA NICs that do not support RDMA. We configure NCCL to use multiple TCP connections that achieve the best performance. With PPG that gives low network RTTs between VMs, UCCL achieves up to 4.1/2.3 higher performance than NCCL for small/large message ranges; without PPG, UCCL achieves up to 2.7/2.1 higher performance. This huge gain is because UCCL implements an efficient multipath transport atop the fast user-space packet IO technique AF_XDP, thus saving significant user-kernel context switching overhead and heavy-weight networking stack traversing in kernel TCP. For data sizes exceeding 16MB, both AF_XDP and TCP experience bottlenecks due to no GPUDirect support.
C.3. UCCL Scalability
C.3.1. Varying number of CPU cores on EFA
Figure 17 shows how the number of CPU cores per NIC impacts UCCL performance on EFA. As mentioned in §6.4.3, 2 cores are sufficient to saturate the EFA line rate, even for the connection-intensive all-to-all collective. Overall, thanks to chained posting, UCCL over UD is able to use 1 core to handle 100 Gbps unidirectional traffic on EFA NICs.
C.3.2. Varying number of GPUs on EFA
Figure 18 shows how UCCL scales with the number of GPUs on EFA. As expected, with less GPUs, UCCL achieves lower latency and higher bus bandwidth. But eventually, UCCL is able to approach line rate with 64MB data size. UCCL leverages the connection-less UD on EFA NICs, thus not suffering from QP scalability issues.
C.3.3. Varying number of paths on EFA
Figure 19 varies the number of paths on the EFA testbed (which crosses racks with multiple network paths). The high-level take here is that multipathing helps mitigate network congestion caused by, e.g., flow collisions. We expect multipath transport to shine more on a larger testbed with more network paths.
C.4. Design Drill-Down
C.4.1. Impact of chunk size and LB policy
This experiment aims to study the impact of chunk size and LB policy on the performance of a broad range of ML transports (not just the ones implemented in UCCL). To this end, we use a packet-level network simulator htsim (Consortium, 2023) and implement UEC standard multipath transports based on their papers (Le et al., 2024; Bonato et al., 2024; Olteanu et al., 2022). We further modify the transport implementation in htsim to vary chunk size and LB policy, e.g., based on ECN or RTT, connection splitting or not. We vary the chunk size by changing the MTU size in the simulator without modifying any transport simulation code. We vary the LB policy by maintaining per-path ECN/RTT and modifying the code where the transport sender selects the path for each packet. Similar to prior work (Raiciu et al., 2011; Cao et al., 2013; Le et al., 2024; Bonato et al., 2024), we focus on the permutation traffic pattern to stress-test the transport, where each NIC streams traffic to another and no one receives more than one stream. We simulate 1024 400G NICs under a fully-provisioned three-tier fattree, and each NIC uses 256 paths to stream 64MB traffic with an ideal completion time of 1.28ms.
Table 4 shows the permutation traffic completion time under different designs. Overall, using 32KB chunk size and switching to RTT degrade transport performance by 17.9% for sender-driven transport and only 2.8% for receiver-driven, while connection splitting does not degrade. With a 16KB chunk size, the performance degradation of sender-driven transport becomes only 4.1%. Receiver-driven performs better because EQDS leverages in-switch packet trimming to quickly detect network congestion and react (Handley et al., 2017; Olteanu et al., 2022). Overall, this experiment shows that control coalescing indeed causes transport performance degradation, but the degradation is moderate in most cases; meanwhile, connection splitting has a negligible impact on transport performance.
Permulation traffic completion time (ms) | Sender-driven | Receiver-driven |
4KB + ECN (UEC default) | 1.45 | 1.43 |
32KB + ECN | 1.62 (+11.7%) | 1.47 (+2.8%) |
32KB + RTT | 1.71 (+17.9%) | 1.47 (+2.8%) |
32KB + RTT + ConnSplit | 1.70 (+17.2%) | 1.47 (+2.8%) |
16KB + RTT + ConnSplit | 1.51 (+4.1%) | 1.42 (-0.7%) |
C.4.2. Impact of kernel fusion.
This experiment aims to quantify the overhead of scattered memcpy and justify UCCL’s design choice of employing kernel fusion over kernel launching (§3.1). In the kernel launching approach, UCCL launches a dedicated copy thread (in CPU) for each UCCL engine at boot time; after receiving all packets of a transport buffer, the engine notifies its associated copy thread via a shared-memory queue; then the copy thread launches a GPU kernel to perform scattered memcpy asynchronously (this kernel is different from the reduction kernel). We further optimize the kernel launching overhead with adaptive batching (Belay et al., 2014). Note that we cannot run the GPU kernel persistently, as this would cause deadlock with the commonly-used cudaDeviceSynchronize(). We also emulate the performance with no scattered memcpy by skipping packet reassembly entirely and disabling data correctness checks in NCCL-tests.
Figure 20 shows the collective performance. Scattered memcpy introduces minor performance overhead compared to no memcpy (less than 8% and 5% for allreduce and all-to-all, respectively). Kernel launching performs worse than kernel fusion for both small message latency and large message bandwidth. The suboptimal performance of Kernel launching is because of the high kernel launching overhead, especially for small messages.
C.4.3. PCIe overhead
This experiment aims to study the PCIe overhead introduced by UCCL’s multipath design. We measure the PCIe overhead on our CX testbed, while the AWS virtualization layer prevents us from accessing low-level PCIe metrics. We rerun the all-to-all experiments in §6.1 and quantify MMIO events and CPU-NIC PCIe traffic using pcm-pcie (Intel, 2025). Figure 21 and 22 show the results. As expected, UCCL incurs higher MMIO activity and PCIe bandwidth consumption than the vanilla NCCL atop CX-7. UCCL incurs more MMIO events because it posts more verbs: UCCL transfers a small 32KB chunk per verb, while the vanilla NCCL atop CX-7 transfers a transport-buffer-sized message (e.g., default 128KB) per verb. The extra PCIe traffic of UCCL primarily comes from NIC swapping QPs (i.e., fetching uncached QP contexts from CPU memory) as UCCL uses 256 UC/RC QPs for multipath. UCCL UC consumes slightly more bandwidth than RC because it implements reliability and selective retransmission in software on the CPU.
We conclude that the number of MMIO events and extra PCIe traffic increase significantly with smaller chunk sizes, but the number of QPs does not nearly affect them. UCCL makes a reasonable trade-off between the control decision granularity and performance based on this observation and uses a 32KB chunk size as default (§3.3). We note that the extra PCIe bandwidth overhead is minor compared to the PCIe link capacity. For example, the capacity of PCIe 5.0 16 on CX is 512 Gbps in each direction, and the overhead of UCCL (32KB, 60k QPs) accounts for less than 2.5% (12/512=2.3%).

### v2

@@
 This software extensibility brings in transport innovations that cannot be achieved in hardware for ML workloads, e.g., a multipath transport to resolve flow collisions.
-ML collectives atop UCCL achieve up to 3.3 higher performance compared to an industry solution.
+ML collectives atop UCCL achieve up to 4.5 higher performance compared to existing RDMA NICs.
 1. Introduction
@@
 To address the second challenge, UCCL leverages an array of techniques tailored to the unique characteristics of ML workloads.
-On the one hand, UCCL leverages GPUDirect (Corporation, 2025d) to alleviate CPU overhead.
+On the one hand, UCCL leverages GPUDirect (Corporation, 2025f) to alleviate CPU overhead.
 On the other hand, UCCL employs control coalescing to make transport control decisions for every 32KB data chunk rather than every packet; this works well as many transport decisions such as CC do not require per-packet reaction, but just per-RTT to avoid overreaction (Li et al., 2019; Google, 2024).
@@
 UCCL leverages 256 RDMA QPs (Queue Pairs) per connection to perform multipathing, without worrying about the high QP context swapping overhead highlighted in previous work (Wang et al., 2023); this is because ML workloads feature bulk data transfer with mostly MTU-sized large packets, which effectively amortizes the swapping overhead.
-UCCL’s software transport is practical and economical in modern GPU servers, which usually have hundreds of powerful CPU cores (in fact, these CPU cores are often heavily underutilized).
+UCCL’s software transport is practical and economical in modern GPU servers, which usually have hundreds of powerful CPU cores; in fact, these CPU cores are often heavily underutilized.
+For example, from our private conversation with the model training team of a major GPU vendor, their CPU utilization is averagely 14.5% out of 128 cores, when using an internal version of Megatron-LM (NVIDIA Corporation, 2025a) for model training.
 UCCL provides an expressive yet easy-to-use interface.
@@
 First, we implement a multipath transport protocol that mitigates flow collisions by leveraging packet spraying, i.e., randomly sending packets from a single connection across different paths (Dixit et al., 2013).
-This transport implementation achieves 3.3 higher throughput for collective communication over AWS’ existing solution, SRD (Shalev et al., 2020) on EFA NICs, which translates into up to 1.4 speedup for two ML applications.
+Compared to existing RDMA transport on hardware, ML collectives using our transport achieve up to 4.5 higher throughput on NVIDIA ConnectX-7 NICs, and up to 1.9 higher on Broadcom Thor-2 NICs with rail-optimized topology.
 Second, we implement a receiver-driven protocol EQDS (Olteanu et al., 2022) to handle network incast in MoE-like workloads, achieving 4.9 better message tail latency over InfiniBand built-in transport.
@@
 These case studies highlight that UCCL can effectively enable innovations in the transport protocols that are hard to implement in today’s network stack without expensive and time-consuming changes.
+UCCL is open-sourced at https://github.com/uccl-project/uccl.
 2. Background and Motivation
@@
 GPU networking is extremely heterogeneous.
-At the high level, GPU collective communication libraries like NVIDIA NCCL (NVIDIA Corporation, 2025) and AMD RCCL (AMD, 2025) use RDMA and kernel TCP (non-RDMA) for inter-server networking, with RDMA preferred since it is faster and more efficient.
+At the high level, GPU collective communication libraries like NVIDIA NCCL (NVIDIA Corporation, 2025b) and AMD RCCL (AMD, 2025) use RDMA and kernel TCP (non-RDMA) for inter-server networking, with RDMA preferred since it is faster and more efficient.
 RDMA provides various communication primitives called Queue Pairs (QPs) including Reliable Connection (RC), Unreliable Connection (UC), and Unreliable Datagram (UD):
@@
 Once the ML application calls a collective like allreduce, the collective library launches a reduction kernel on each participant GPU to handle data reduction/copy.
-Next, the sender CPU issues an RDMA write over RC QPs to transfer data.
-The receiver CPU polls a completion flag in its memory, which the sender sets upon write completion.
+Next, the sender CPU issues multiple RDMA writes over RC QPs to transfer data chunk by chunk.
+The receiver CPU polls completion flags in its memory, which the sender sets upon write completions.
 The library manages a set of transport buffers on the GPU memory to buffer the RDMA data, and relies on the GPU kernel to copy data between the transport buffers and application tensor buffers.
@@
 Recent MoE serving workloads are prone to network incast problems.
-In DeepSeek’s online deployment of their V3 model (Liu et al., 2024), each of the 320 GPUs holds a single expert module, where hidden states are exchanged between expert modules across GPUs, i.e., Expert Parallelism (EP).
+In DeepSeek’s online deployment of their 671B V3 model (Liu et al., 2024), each of the 320 GPUs holds a single expert module, where hidden states are exchanged between expert modules across GPUs, i.e., Expert Parallelism (EP).
 As the request pattern and load change over time, some experts become much hotter than others, receiving more network traffic from other experts leading to network incast issues.
 DeepSeek reports the hottest expert could receive 10 more load than the average one.
-This network incast can be better handled by receiver-driven CC (Handley et al., 2017; Olteanu et al., 2022) which controls last-hop congestion—unfortunately, there is no receiver-driven CC on commercial RDMA NICs.
+Expert load balancing algorithms such as EPLB (DeepSeek AI, 2025) try to balance the load by dynamically replicating experts.
+However, this happens at a much slower pace (e.g., 10 minutes (DeepSeek, 2025)) to avoid the high cost of moving experts, thus unable to handle transient incast.
+Such transient network incast can be better handled by receiver-driven CC (Handley et al., 2017; Olteanu et al., 2022), which controls last-hop congestion—unfortunately, there is no receiver-driven CC on commercial RDMA NICs.
 Application-transport codesigns.
@@
 Datacenters usually consist of multiple generations and vendors of RDMA NICs due to continuous expansion, cost optimization, and to avoid vendor lock-in.
-While NVIDIA, Broadcom, and AMD all have 400 Gbps RDMA NICs for ML (NVIDIA Corporation, 2024a; Broadcom, 2024; AMD Pensando, 2024), they come with subtly different control path logic like packet reliability and CC; this heterogeneity reduces achievable bandwidth by 2-33 when communicating between NICs from different generations/vendors, as reported by Alibaba (Li et al., 2023).
-Instead, if we can extensibly align these NICs’ control path logic in software, we could avoid such a severe performance drop.
+While NVIDIA, Broadcom, AMD, and more vendors all have 400 Gbps RDMA NICs for ML (NVIDIA Corporation, 2024a; Broadcom, 2024; AMD Pensando, 2024), they come with subtly different control path logic such as packet reliability and CC.
+In practice, this heterogeneity reduces achievable bandwidth by 2-33 when communicating between NICs from different generations/vendors, as reported by Alibaba (Li et al., 2023).
+Prior work Flor (Li et al., 2023) has shown that extensibly aligning these NICs’ control path logic in software could avoid such a severe performance drop.
 2.3. Prior Work on Extensibility
@@
 AMD Pensando SmartNICs (AMD Pensando, 2024) support using P4 language (Bosshart et al., 2014) to program their transport layer, but P4 has limited programmability, e.g., hard to implement efficient loss recovery; it is also unclear if they could support receiver-driven CC.
-AWS EFA SmartNICs (Services, 2025) implement a proprietary multipath reliable transport SRD (Shalev et al., 2020) using NIC ARM cores.
-However, they perform poorly on collectives, especially for the connection-intensive all-to-all (see §6.1).
-We attribute this to the fundamentally limited processing power and cache capacity of these wimpy RISC cores (Phothilimthana et al., 2018; Liu et al., 2019).
 FPGA-based SmartNICs provide higher performance but have limited extensibility due to hardware resource constraints (Wang et al., 2023).
+AWS EFA SmartNICs (Services, 2025) implement a proprietary multipath reliable transport SRD (Shalev et al., 2020) with out-of-order packet delivery using NIC ARM cores, to address network congestion in HPC and ML workloads.
+The SRD protocol is implemented using EFA-specific firmware and supports live upgrade with good extensibility.
+However, we empirically find that the EFA SmartNICs on AWS p4d.24xlarge GPU VMs perform poorly for connection-intensive all-to-all collectives (see §6.1).
+We attribute this to the limited processing power and cache capacity of SmartNIC ARM cores due to power constraints, as also demonstrated by prior research on publicly available SmartNICs (Phothilimthana et al., 2018; Liu et al., 2019; Schuh et al., 2021).
+Note that the above all-to-all measurement uses the AWS EFA NICs on the p4d.24xlarge instances, and thus it may not apply to their new-generation EFA NICs on p5/p5en/p6 instances.
+Specifically, the ever-changing EFA firmware and upgraded EFA hardware may lead to different performance results for connection-intensive all-to-all collectives.
+In general, if these SmartNICs get upgraded with higher processing power and cache capacity to better handle all-to-all, we view it as an echo to UCCL’s high-level methodology of software extensibility for GPU networking.
 Leveraging CPUs.
@@
 Overall, these protocols leverage various congestion signals such as Explicit Congestion Notification (ECN) (Alizadeh et al., 2010), RTT (Mittal et al., 2015), and packet trimming status (Cheng et al., 2014; Handley et al., 2017; Olteanu et al., 2022) to make multipath CC and LB decisions.
-How to separate control and data (§3.1)? | How to reassemble packets on receivers (§3.1)? | How to enable multipathing (§3.2)? | How to build efficient software transport (§3.3)? | |
-RC, UC | Write with immediate | RDMA NIC handles | Different QP connections | Control coalescing |
-UD | Send/Recv with scatter-gather | GPU does scattered memcpy | Src-dst QP combinations | Chained posting |
-AF_XDP | – | CPU handles | Different UDP ports | User-space packet IO |
 3. UCCL Design
@@
 UCCL also supports extending transport for non-RDMA NICs through the AF_XDP user-space packet IO, bypassing conventional kernel TCP stacks.
-Table 1 summarizes the UCCL techniques to achieve transport extensibility across different communication primitives.
 We now describe them in the rest of this section.
@@
 UCCL chooses UC as the preferred QP whenever available on the RDMA NIC, similar to Flor (Li et al., 2023), because it supports efficient segmentation and reassembly offloading to the NIC (in contrast to UD), while bypassing hardware-baked CC, loss recovery, and out-of-order packet handling (in contrast to RC).
-Under UC, UCCL uses the efficient RDMA verb of write with immediate to transfer data chunks; this verb operates in a two-sided mode, so that both sender and receiver CPUs can react to the data transfers.
+As shown in Figure 3, UCCL uses the efficient RDMA verb of write with immediate to transfer data chunks over UC; this verb operates in a two-sided mode, so that both sender and receiver CPUs can react to the data transfers.
 For the data path, the sender CPU specifies the addresses of source and destination data chunks when issuing the verb, and then the sender NIC will automatically segment the source data chunk into MTU-sized packets with packet headers prepended and send them out.
@@
 For the control path, write with immediate allows carrying a 32-bit imm_data from the sender CPU to the receiver CPU, serving as the control header for UCCL transport.
+Figure 3 shows an example.
 UC guarantees that any successfully arrived chunk will generate a CQE with the imm_data embedded, which is then consumed by the receiver CPU.
@@
 One of the key motivations for GPU network extensibility is to harness the multipath capacity of modern datacenter networks (§2.2).
-UCCL achieves this by using multiple UC, RC, or UD QPs.
-Basically, network traffic from different QPs will likely go through different network paths, as both RoCE and Infiniband usually use ECMP (Equal-Cost Multi-Path) for multipath routing with source and destination QP numbers as the hash inputs (Gangidi et al., 2024; NVIDIA Corporation, 2024b).
+UCCL achieves this by using multiple UC, RC, or UD QPs, as shown in Figure 5.
+Basically, network traffic from different QPs will likely go through different network paths, as both RoCE and Infiniband usually use ECMP (Equal-Cost Multi-Path) for multipath routing with source and destination QP numbers as the hash inputs (Gangidi et al., 2024; NVIDIA Corporation, 2024c).
 For UC and RC, UCCL by default uses 256 QPs, which provides maximum 256 different network paths as used by recent transport research (Le et al., 2024; Bonato et al., 2024).
@@
 This bandwidth drop is caused by the QP swapping overhead: the NIC can only hold/cache limited QP contexts on its SRAM, and must spill/swap the excessive rest to the host DRAM over PCIe, incurring frequent QP swapping.
-Surprisingly, we do not observe such a severe performance drop for collective communication—we only see 17% drop when scaling RC QPs to 60k (see §6.4.1); for UC with a smaller QP context size, the drop is even negligible.
+Surprisingly, we do not observe such a severe performance drop for collective communication—Figure 6 shows only 17% drop when scaling RC QPs from 60 to 60k, and a negligible drop for UC (that has a smaller QP context size than RC).
 There are two reasons behind this counter-intuitive phenomenon.
@@
 The next question is how to efficiently implement a software multipath transport to support the high bandwidth in GPU networking.
-This is challenging as a single GPU server could have 8400 Gbps RDMA NICs, totaling 3.2 Tbps bandwidth bidirectionally (Huang et al., 2012); the next generation RDMA NIC will achieve 800 Gbps (Corporation, 2025c), rendering 6.4 Tbps bandwidth.
+This is challenging as a single GPU server could have 8400 Gbps RDMA NICs, totaling 3.2 Tbps bandwidth bidirectionally (Huang et al., 2012); the next generation RDMA NIC will achieve 800 Gbps (Corporation, 2025e), rendering 6.4 Tbps bandwidth.
 As a reference, Google’s software transport Snap (Marty et al., 2019) can handle 80 Gbps traffic on a CPU core (though they do not use RDMA NICs).
-Our goal is to use 1 CPU core to handle 400G unidirectional traffic.
+Our goal is to use 1 CPU core to handle 400G unidirectional traffic (i.e., 2 cores for 400G bidirectional traffic; excluding possible pacer cores for receiver-driven CC).
 To this end, we leverage the following techniques:
@@
 Concretely, the WQEs of these 32 verbs are chained together through the next pointer in previous WQEs, and get posted to the RDMA NIC in one MMIO write.
-Small message bypassing multipath.
-Using 256 UC/RC QPs for small messages still incurs a high performance penalty in terms of latency; this is because these small message transfers cannot fully amortize the QP swapping overhead.
-To optimize this, UCCL empirically bypasses multipathing for messages under 8KB and transmits them directly via a dedicated UC/RC QP.
-We argue this bypassing design for small messages does not impact the overall benefits of multipathing, which usually begins to matter at high traffic load with large message transfers (Le et al., 2024).
 3.4. Congestion Signals
@@
 With more efficient loss recovery in UCCL, ML workloads could possibly run in lossy datacenter networks without PFC (§2.2).
+Name | # of Servers | Network | Topology | MTU | NIC | GPU | CPU |
+CX_IB | 2 | InfiniBand | Same rack | 4KB | NVIDIA ConnectX-7 400G 8 | NVIDIA H100-80G 8 | 128 cores |
+CX_ETH | 6 | Ethernet | Cross racks, fat-tree | 4KB | NVIDIA ConnectX-7 400G 8 | NVIDIA H100-80G 8 | 160 cores |
+AMD | 4 | Ethernet | Cross racks, rail-optimized | 4KB | Broadcom Thor‑2 400G 7 | AMD MI300X-192G 7 | 128 cores |
+EFA | 4 | Ethernet | Cross racks, fat-tree | 9KB | AWS EFA 100G 4 | NVIDIA A100-40G 8 | 96 cores |
 We note that the latest generation of NVIDIA RDMA NICs (NVIDIA Corporation, 2024a) has implemented a limited form of selective retransmission with a small fixed tracking window (to keep low SRAM usage).
@@
 5. Implementation
-We implement UCCL in 27.4K lines of C++ as a network plugin to NCCL, the standard collective library for NVIDIA GPUs.
-Our current implementation supports NVIDIA RDMA NICs via RC and UC, AWS EFA NICs via UD, and AWS ENA non-RDMA NICs via AF_XDP.
-UCCL leverages the popular libibverbs for RDMA NICs and kernel built-in AF_XDP for non-RDMA NICs; therefore, it should naturally support other vendors’ NICs.
-To enable scattered memcpy when using UD, we add or modify 170 LoC to the NCCL codebase, including a scattered memcpy GPU kernel and C++ code that passes packet pointers to the kernel via CPU-GPU shared memory (created by cudaHostAlloc()).
+We implement UCCL in 28.4K lines of C++ as a network plugin to NCCL/RCCL, the standard collective library for NVIDIA/AMD GPUs.
+Our current implementation supports NVIDIA RDMA NICs via RC and UC, Broadcom RDMA NICs via RC, AWS EFA NICs via UD/SRD, and AWS ENA non-RDMA NICs via AF_XDP.
+UCCL leverages the standard libibverbs for RDMA NICs and kernel built-in AF_XDP for non-RDMA NICs; therefore, it should naturally support other vendors’ NICs.
+To enable scattered memcpy when using UD for AWS EFA NICs, we add or modify 170 LoC to the NCCL codebase, including a scattered memcpy GPU kernel and C++ code that passes packet pointers to the kernel via CPU-GPU shared memory (created by cudaHostAlloc()).
 We added two new interfaces to the NCCL network plugin system: irecv_scattered that receives data in a list of scattered packets, and irecv_free_ptrs that frees the scattered packet buffers after scattered memcpy.
 UCCL also supports running NCCL with multiple processes.
-Name | # of Servers | Network | Topology | NIC | GPU | CPU |
-CX | 2 | InfiniBand | Same rack | NVIDIA ConnectX-7 400G 8 | H100-80G 8 | Intel Xeon Platinum 8462Y+ (128 cores) |
-EFA | 4 | Ethernet | Fat-tree cross racks | AWS EFA 100G 4 | A100-40G 8 | Intel Xeon Platinum 8575CL (96 cores) |
 There are some implementation details worthy of note.
@@
 Table 1 describes our evaluation testbeds.
-The CX testbed under the same rack mainly serves as a performance stress-testing for UCCL software implementation, as it has no network congestion or flow collisions.
-We failed to rent a larger testbed from public cloud providers because they have strict GPU quota limits for individuals.
-Instead, we try to simulate a larger testbed by disabling NCCL NVLink and SHM communication for some tests.
-This forces GPUs within a server to communicate through the network, and each GPU essentially behaves like a virtual server, e.g., CX becomes 16 virtual GPU servers each with 400G bandwidth.
+The CX_IB testbed under the same rack mainly serves as a performance stress-testing for UCCL software implementation, as it has no network congestion or flow collisions.
+The CX_ETH testbed across two racks evaluates how UCCL handles network congestion and flow collisions.
+The AMD and EFA testbed across racks evaluate the genericity of UCCL (i.e., applying to AMD GPUs and Broadcom/EFA NICs).
+Each AMD server has one NIC down out of 8; therefore, we only use 7 GPUs for evaluation.
+For some tests, we try to simulate a larger testbed by disabling NCCL NVLink (or RCCL xGMI) and SHM communication.
+This forces GPUs within a server to communicate through the network, and each GPU essentially behaves like a virtual server, e.g., CX_IB becomes 16 virtual GPU servers each with 400G bandwidth.
 Experiment setup.
-For collective performance evaluation, we use NCCL v2.23.4 (Corporation, 2024b) and NCCL-tests 9d26b84 (Corporation, 2024a), the latest versions when we start integrating UCCL into NCCL, and focus on the representative allreduce and all-to-all collectives.
-NCCL-tests vary the collective message size and measure the achieved latency and bus bandwidth; the latency is inversely proportional to the bandwidth.
+For collective performance evaluation, we use NCCL v2.23.4 (Corporation, 2025d) and NCCL-tests 9d26b84 (Corporation, 2025c), the latest versions when we start integrating UCCL into NCCL, and focus on the representative allreduce and all-to-all collectives.
+NCCL-tests vary the collective message size and measure the achieved bus bandwidth.
+We focus on message sizes from 1MB to 1GB, commonly used in real ML workloads (Qian et al., 2024; Wang et al., 2025; Xu et al., 2025).
+For AMD GPUs, we use RCCL 532f54c (Advanced Micro Devices, 2025b) and RCCL-test 5b27b96 (Advanced Micro Devices, 2025a).
 Comparison baselines.
-On the CX testbed, we compare UCCL to the NCCL built-in RDMA support that uses CX-7 RC.
+On the CX_IB testbed, we compare UCCL to the NCCL built-in RDMA support that uses CX-7 RC.
 UCCL uses 256 UC/RC QPs for each NIC pair, while the NCCL built-in uses QP scaling with 4 RC QPs per connection (Gangidi et al., 2024; Le et al., 2024).
-UCCL uses CUBIC CC on CX unless specified, which performs better than Swift in the lossless InfiniBand.
+UCCL uses CUBIC CC on CX_IB unless specified, which performs better than Swift in the lossless InfiniBand.
 Since InfiniBand has no packet loss for CUBIC to reduce cwnd, to avoid severe network congestion, we enforce a maximum cwnd value in UCCL CUBIC CC to limit the maximum inflight bytes, similar to TCP flow control.
@@
 As the cost of extensible software transport, NCCL atop UCCL uses more CPU cores than vanilla NCCL.
-By default, the vanilla NCCL uses 2 cores per GPU, while UCCL uses 2 more cores per NIC to run the engine threads (3 more cores per NIC for receiver-driven CC due to the pacer).
-For example, on CX, they use 16 and 32 cores respectively; on EFA, 16 and 24 respectively.
+By default, the vanilla NCCL uses 2 cores per GPU, while UCCL uses 2 more cores per NIC to run the engine threads, with 1 additional core per NIC for receiver-driven CC due to the pacer.
 6.1. Collective Performance
-On CX.
+On CX_IB.
 Figure 7 compares the allreduce and all-to-all performance of UCCL vs.
-ConnextX-7 on the CX testbed.
+ConnextX-7 on the CX_IB testbed.
 UCCL UC/RC performs almost the same as ConnectX-7 with various data sizes, except that UCCL UC performs slightly worse (i.e., less than 4%) than ConnextX-7 and UCCL RC for allreduce after 128MB.
 For allreduce, this exception is because of the extra overhead of handling packet reliability in UC; for all-to-all, ConnextX-7 and UCCL RC do not perform better than UCCL UC, because all-to-all is connection-intensive, causing more QP swapping for RC.
-UCCL exhibits lower latency for small messages than ConnectX-7 since UCCL adaptively bypasses multipath (§3.3), while ConnectX-7 always uses 4 RC QPs per connection.
 These experiments confirm that UCCL can make highly efficient control decisions in software, reaching the performance of ASIC-based RDMA NICs (i.e., NVIDIA ConnectX-7) for ML collectives.
+On CX_ETH.
+Figure 8 compares the performance of UCCL vs.
+ConnextX-7 on the CX_ETH testbed.
+For NCCL over CX-7, we use multiple QPs for each connection between two GPUs by adjusting NCCL_IB_QPS_PER_CONNECTION.
+However, it still suffers from significant performance drops under larger message sizes and lower peak performance.
+This is because flow collisions in the fat-tree topology cause severe network congestion, which in turn leads to exponential backoffs in the CX-7 CC mechanism.
+Instead, UCCL is able to scale up performance stably with message size increasing, by making smarter LB and CC decisions in software.
+Overall, UCCL outperforms CX-7 by up to 2.32/1.60/1.24 (corresponding to 4/8/16 QPs) and 1.79/3.82/4.54 for allreduce and all-to-all, respectively.
+On AMD.
+Figure 9 shows the performance comparison on the AMD testbed.
+This testbed has a rail-optimized topology that helps reduce network congestion when used together with the PXN technique (NVIDIA Corporation, 2022).
+Even under this topology, UCCL still outperforms Thor-2 by up to 1.34/1.23/1.08 (corresponding to 4/8/16 QPs) and 1.62/1.62/1.93 for allreduce and all-to-all, by better controlling network congestion.
 On EFA.
 Figure 10 compares the collective performance of UCCL and SRD on the EFA testbed.
-UCCL CUBIC and EQDS achieve similar performance with significant improvement over SRD (except for 256MB allreduce): for allreduce, UCCL outperforms SRD by up to 1.37 and 1.27 for small and large message ranges; for all-to-all, the speedup is up to 2.93 and 3.27 respectively.
-UCCL outperforms SRD because beefy CPU cores are faster in making transport decisions than wimpy ARM cores on EFA NICs (that run SRD), especially when handling connection-intensive all-to-all.
+UCCL CUBIC and EQDS achieve similar performance with significant improvement over SRD (except for 256MB allreduce): for allreduce, UCCL outperforms SRD by up to 1.27; for all-to-all, the speedup is up to 3.27.
+UCCL outperforms SRD because beefy CPU cores are faster in making transport decisions than wimpy ARM cores on p4d.24xlarge EFA NICs (that run SRD), especially when handling connection-intensive all-to-all.
 Surprisingly, simple CUBIC CC performs very well.
@@
 The small performance gap after 256MB is caused by the extra control header added by UCCL, while SRD directly reuses the UDP header as its control header (UCCL cannot reuse as it is not exposed to CPU software).
+UCCL could also directly leverage the SRD protocol of EFA NICs to send/recv individual datagram packets, denoted as UCCL SRD.
+UCCL SRD does connection management and multipath load balancing in CPU, and packet reassembly in GPU (§3.1), overcoming the limited processing power and cache capacity of SmartNICs on p4d.24xlarge.
+As a result, it achieves similar high performance to UCCL CUBIC and EQDS.
 Figure 11 shows the collective performance comparison on the EFA testbed when NVLink and SHM are enabled (meaning a smaller testbed).
 Even though a major amount of data traffic goes through the high-bandwidth NVLink, UCCL still achieves much higher or comparable collective performance than SRD, i.e., up to 1.57 for allreduce and 2.14 for all-to-all.
-We note that all-to-all with NVLink on performs worse than NVLink off (Figure 6).
-This is because our current UCCL implementation does not support the NCCL PXN (NVIDIA Corporation, 2022) optimization when NVLink is on.
-Basically, PXN aggregates network messages destined to the same server, and uses the destination NVLink to segment them into different GPUs, which vastly improves small message latency.
-However, it involves a complicated multireceive mechanism that causes engineering difficulties for UCCL UD integration (that requires changing the NCCL codebase), which we leave as future development.
-Nevertheless, UCCL still performs better than SRD under small messages in this experiment.
 6.2. Application Performance
+UCCL is motivated by the practical needs of both LLM training and serving workloads.
 We run two applications to evaluate how UCCL improves ML workload performance on the EFA testbed.
@@
 Another is a DeepSeek-V3-like MoE serving application.
-However, we fail to directly serve DeepSeek-V3 on our testbed, because (1) it uses a customized communication library rather than the standard NCCL, so it cannot directly use UCCL, and (2) our testbed does not have enough GPU memory to fit all model weights.
+However, directly serving DeepSeek-V3 on our testbed is impractical, because (1) it uses a customized communication library rather than the standard NCCL, so it cannot directly use UCCL, and (2) it requires a large number of GPUs to run meaningful expert parallelism, e.g., 320 GPUs by DeepSeek (Liu et al., 2024) and 96 GPUs by SGLang (The SGLang Team, 2025).
 Instead, we get a realistic DeepSeek-V3 trace including GPU compute time and network message sizes (i.e., hidden state sizes) from (zartbot/shallowsim, 2025), and emulate the computation and communication behaviors with PyTorch and NCCL.
@@
 Such network incast can be caused by the Expert Parallelism in MoE serving (§2.2) and gather/reduce collectives in multi-level parallelism of LLM training (Barham et al., 2022; Grattafiori et al., 2024).
-This experiment tries to create a scenario where network incast co-exists with other collective traffic, and evaluates how UCCL performs compared to the RDMA hardware transport on CX.
+This experiment tries to create a scenario where network incast co-exists with other collective traffic, and evaluates how UCCL performs compared to the RDMA hardware transport on CX_IB.
 We co-locate 15-to-1 incast traffic and 16-NIC permutation traffic (Raiciu et al., 2011; Cao et al., 2013; Le et al., 2024; Bonato et al., 2024), where each NIC streams data to another and no NIC receives more than one stream, representing typical collectives.
@@
 the sender-driven InfiniBand CC in ConnectX-7.
-Figure 9 and 9 show the FCT distributions for incast traffic and permutation traffic, respectively.
+Figure 13 and 13 show the FCT distributions for incast traffic and permutation traffic, respectively.
 Compared to InfiniBand, UCCL EQDS reduces P99/P99.9 latency by 1.73/1.72 for incast traffic and 4.50/4.88 for permutation traffic.
@@
 To this end, we instrument different packet loss rates in software when running NCCL collectives over UCCL UC.
-Unfortunately, we cannot instrument packet loss for our CX NICs as that would require reconfiguring the NIC or switch behaviors, both of which we do not have access to.
-Instead, we cite comparable numbers from prior literature.
+Our packet loss instrumentation has considered the chunking nature in UCCL, where any packet loss will cause the whole chunk loss.
+Unfortunately, we cannot instrument packet loss for our CX_IB NICs as that would require reconfiguring the NIC or switch behaviors, both of which we do not have access to.
+Instead, we cite comparable numbers from prior literature Flor (Li et al., 2023).
+We pick two GPUs from the AMD testbed, each using one NIC for communication.
+The client NIC establishes a connection with the server NIC and keeps 16 inflight messages.
+Similar to Flor (Li et al., 2023), we use a single QP for the connection, and disable congestion control to avoid sending rate backoff.
+We vary the message size (32KB1MB) and measure the goodput under different packet drop ratios.
 Figure 14 shows the results.
-UCCL has only 3-23%/22-41% performance drop for allreduce and all-to-all under 0.01% packet loss rate, compared to the reported 77% drop of RDMA hardware transport in Flor (Li et al., 2023) (combining Flor Figure 6(a) and 7(a) under 1/16384 loss rate).
-With 1% high packet loss rate, UCCL still retains 58%/72% of the performance for 1GB messages, compared to old generation RDMA NICs that achieve near-zero throughput as reported in DCQCN (Zhu et al., 2015) (Figure 21).
-This confirms the effectiveness and advantage of UCCL loss recovery despite control coalescing.
+UCCL has only 1% performance drop under 1/16384, 1/4096 drop ratio, compared to the reported 26%42% drop of RDMA hardware transport in Flor (Li et al., 2023, Figure 7).
+Even under a high drop ratio of 1/1024 and 1/256, UCCL only experiences a performance drop of 6%30%, compared to the reported 59%76% drop in RDMA hardware transport.
 6.4. UCCL Scalabality
-6.4.1. Varying number of QPs
-We vary the number of QPs per NIC pair from 256 to 4096; given that 1 NIC connects to 15 other NICs in all-to-all, this results in 3840 (3.75K) to 61440 (60K) QPs per NIC.
-With 60K QPs per NIC, we essentially model the QP swapping overhead for UCCL all-to-all across 241 GPUs using 256 paths per connection (i.e., 61440/256=240, plus 1 to include itself) or 961 GPUs using 64 paths.
-This has covered the largest-scale collective as far as we know, i.e., 256 GPUs reported by Meta (Gangidi et al., 2024) (DeepSeek-V3 limits each token to be dispatched to 4 servers, though with 320 GPUs in total (Liu et al., 2024)).
-Figure 11 shows the results.
-For RC, using 60K QPs incurs a maximum 17% performance drop compared to 3.75K QPs (that have similar performance as the vanilla NCCL).
-For UC, the performance remains nearly unchanged across different numbers of QPs.
-As mentioned in §3.2, this is because UCCL’s chunked data transfers (i.e., control coalescing) for large ML messages effectively amortize the QP swapping overhead while GPUDirect avoids PCIe traffic contention.
-Note that UCCL on EFA does not have QP scalability issues, as UCCL uses connection-less UD there.
 6.4.1. Varying chunk size
-Figure 12 shows the all-to-all performance on the CX testbed when varying chunk sizes.
-Small message performance mostly remains unaffected by different chunk sizes, thus omitted here.
-Our results reveal that large messages do require a larger chunk size to achieve line rate, i.e., 16KB or 4 MTU-sized chunks.
+Figure 15 shows the all-to-all performance on the CX_IB testbed when varying chunk sizes.
+Our results reveal that saturating the line rate requires a somewhat medium chunk size, i.e., 16KB or 4 MTU-sized chunks.
 This experiment demonstrates the performance benefits of control coalescing for UCCL software transport.
@@
 Figure 15 shows the all-to-all performance when varying the number of UCCL engines/CPU cores per NIC.
-Small message latency exhibits negligible difference across different CPU cores, thus omitted here.
 For ASIC-based NICs with segmentation and reassembly offloading, even though one CPU core can only saturate bidirectional 29GB/s, two CPU cores are enough to saturate the full 50GB/s line rate.
-This experiment confirms the high efficiency of UCCL software transport, i.e., 1 CPU core saturating 400G unidirectional traffic.
+The slight performance drop for 16MB when switching from “2CPUs” to “4CPUs” should be caused by runtime variability.
+This experiment confirms the high efficiency of UCCL software transport, i.e., 1 CPU core saturating 400G unidirectional traffic (2 CPU cores for 400G bidirectional).
 We show more scalability studies in Appendix C.3 when varying numbers of CPU cores, GPUs, and paths on EFA.
@@
 Without connection splitting, each UCCL connection is mapped to a single engine out of the two UCCL engines per NIC; with splitting, each UCCL connection dispatches and load balances its messages among the two UCCL engines.
-Our experiment results show that UCCL without connection splitting only achieves 45.7 and 39.9 GB/s max busbw for allreduce and all-to-all on CX respectively, while UCCL with splitting achieves 48.9 and 48.5 GB/s respectively, saturating the line rate.
+Our experiment results show that UCCL without connection splitting only achieves 45.7 and 39.9 GB/s max busbw for allreduce and all-to-all on CX_IB respectively, while UCCL with splitting achieves 48.9 and 48.5 GB/s respectively, saturating the line rate.
 This is because connection splitting enables multiple CPU cores to process inflight messages in a load-balanced manner.
-We anticipate connecting splitting will become more critical as the NIC scales to 800 Gbps (Corporation, 2025c).
+We anticipate connecting splitting will become more critical as the NIC scales to 800 Gbps (Corporation, 2025e).
+Delay () | Allreduce Light | All-to-All Light | Allreduce Heavy | All-to-All Heavy |
+CC decision P50 | 1.7 | 1.7 | 2.1 | 2.9 |
+CC decision P99 | 3.6 | 3.4 | 7.0 | 10.8 |
+ACK turnaround P50 | 2.0 | 3.0 | 4.0 | 7.0 |
+ACK turnaround P99 | 3.0 | 10.0 | 10.0 | 36.0 |
 6.5.2. Congestion signal fidelity in software
@@
 Therefore, this experiment confirms that UCCL software transport has high enough CC fidelity to make precise per-RTT transport decisions (Li et al., 2019; Google, 2024).
-Delay () | Allreduce Light | All-to-All Light | Allreduce Heavy | All-to-All Heavy |
-CC decision P50 | 1.7 | 1.7 | 2.1 | 2.9 |
-CC decision P99 | 3.6 | 3.4 | 7.0 | 10.8 |
-ACK turnaround P50 | 2.0 | 3.0 | 4.0 | 7.0 |
-ACK turnaround P99 | 3.0 | 10.0 | 10.0 | 36.0 |
 We show more design drill-downs in Appendix C.4, including the impact of chunk size, LB policy, kernel fusion (for scattered memcpy), and PCIe overhead analyses.
 7. Discussion
+HW-SW interface.
 UCCL works around many challenges to support extensible transport with efficient multipathing on existing RDMA NICs, at the cost of QP swapping overhead (though not high for ML collectives), control coalescing, and more.
@@
 (2) NIC HW should expose more congestion signals to SW, such as ECN marks and packet trimming information; these signals can be embedded into CQEs just like hardware timestamps.
+GPU-driven communication like DeepEP (DeepSeek, 2025) leverages NVIDIA IBGDA (NVIDIA Corporation, 2025c) to issue RDMA verbs directly from GPUs to the RDMA NICs.
+Although seemingly conflicted with UCCL’s CPU-driven designs, GPU-driven communication could still be made compatible with UCCL.
+The key is to leverage the IBGDA CPU-assisted mode (NVIDIA Corporation, 2024b), where the GPU forwards RDMA requests to a CPU proxy that issues RDMA verbs; with this, we could implement UCCL’s extensible transport layer inside the CPU proxy.
+The trade-off is on performance: NVIDIA has reported that this CPU-assisted IBGDA would sacrifice 10% performance compared to the traditional IBGDA (NVIDIA Corporation, 2024b).
+We intend to leave the integration and performance enhancement of CPU-assisted IBGDA + UCCL as future work.
 8. Other Related Work
@@
 MSCCL (Azure/msccl, 2024) supports customizing collective communication algorithms and could work together with UCCL.
-Lastly, UCCL is inspired by a line of work that targets extensibility for CPU applications, e.g., Google 1RMA (Singhvi et al., 2020), eRPC (Kalia et al., 2019), and RoGUE (Le et al., 2018) for extensible RDMA transport, and SPIN (Bershad et al., 1995), Exokernel (Engler et al., 1995), and VINO (Seltzer et al., 1996) for extensible OS.
+Lastly, UCCL is inspired by a line of work that targets extensibility for CPU applications, e.g., Google 1RMA (Singhvi et al., 2020), eRPC (Kalia et al., 2019), RoGUE (Le et al., 2018), and IO-TCP (Kim et al., 2023) for extensible RDMA transport, and SPIN (Bershad et al., 1995), Exokernel (Engler et al., 1995), and VINO (Seltzer et al., 1996) for extensible OS.
 9. Conclusion
@@
 We hope UCCL opens the door to the productization of new research proposals on network transports for ML workloads.
-We will open source UCCL.
+UCCL is open-sourced at https://github.com/uccl-project/uccl.
 Appendix A Interface to Extensibility
@@
 We evaluate multi-collective performance by setting the environment variable NCCL_TESTS_SPLIT_MASK=0x7 in NCCL-tests.
-Figure 15, 15, 15 and 15 compare the multi-allreduce, multi-all-to-all, multi-allgather and multi-reduce-scatter performance of UCCL and SRD on the EFA testbed.
+Figure 18, 18, 18 and 18 compare the multi-allreduce, multi-all-to-all, multi-allgather and multi-reduce-scatter performance of UCCL and SRD on the EFA testbed.
 UCCL outperforms SRD by up to 1.54, 1.22, 1.46, and 1.44 for multi-allreduce, multi-all-to-all, multi-allgather and multi-reduce-scatter, respectively.
@@
 This experiment aims to study the PCIe overhead introduced by UCCL’s multipath design.
-We measure the PCIe overhead on our CX testbed, while the AWS virtualization layer prevents us from accessing low-level PCIe metrics.
+We measure the PCIe overhead on our CX_IB testbed, while the AWS virtualization layer prevents us from accessing low-level PCIe metrics.
 We rerun the all-to-all experiments in §6.1 and quantify MMIO events and CPU-NIC PCIe traffic using pcm-pcie (Intel, 2025).
-Figure 21 and 22 show the results.
+Figure 24 and 25 show the results.
 As expected, UCCL incurs higher MMIO activity and PCIe bandwidth consumption than the vanilla NCCL atop CX-7.
@@
 We note that the extra PCIe bandwidth overhead is minor compared to the PCIe link capacity.
-For example, the capacity of PCIe 5.0 16 on CX is 512 Gbps in each direction, and the overhead of UCCL (32KB, 60k QPs) accounts for less than 2.5% (12/512=2.3%).
+For example, the capacity of PCIe 5.0 16 on CX_IB is 512 Gbps in each direction, and the overhead of UCCL (32KB, 60k QPs) accounts for less than 2.5% (12/512=2.3%).